### v2

# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

Abstract
[…]
Our analysis of the mixture case hinges on two new covering results. First, we provide a sharp, distribution-free upper bound on the local Hellinger entropy of the class of mixtures of $M$ distributions. Second, we prove an optimal ratio covering theorem for convex sets: for every convex compact set $K\subset\mathbb{R}_{+}^{d}$, there exists a subset $A\subset K$ with at most $2^{O(d)}$ elements such that each element of $K$ is coordinate-wise dominated by an element of $A$ up to a universal constant factor. This geometric result is of independent interest; notably, it yields new cardinality estimates for $\varepsilon$-approximate Pareto sets in multi-objective optimization with convex feasible set.
1 Introduction
We revisit two classical distribution estimation problems. Let $\mathcal{X}$ be a measurable space endowed with a measure $\mu$. In addition, let $p_{1},\ldots,p_{M}$ be a collection of $M\geqslant 2$ known densities on $\mathcal{X}$ with respect to $\mu$, with no assumptions on these densities (in particular, the corresponding distributions need not satisfy any moment or tail conditions, and their supports may differ). We observe an i.i.d. sample $X_{1},\ldots,X_{n}$ drawn from some unknown density $p^{\star}$ for which we only assume either
  1. 1.
    $p^{\star}\in\{p_{1},\ldots,p_{M}\}$, in the Model Aggregation setting;
  2. 2.
    or $p^{\star}\in\operatorname{conv}\{p_{1},\ldots,p_{M}\}$, in the Convex Aggregation (or mixture density estimation) setting.
[…]
In this work, our primary aim is to characterize the best possible high-probability guarantees for both Model Aggregation and Convex Aggregation. More precisely, given $M,n\geqslant 2$ and $\delta\in(0,1/2)$, we aim to identify (up to absolute constant factors) the smallest quantities $\psi_{\mathrm{MA}}(M,n,\delta)$ and $\psi_{\mathrm{CA}}(M,n,\delta)$ such that, for every probability densities $p_{1},\dots,p_{M}$ on a measure space $(\mathcal{X},\mu)$:
  1. 1.
    there exists an estimator $\widehat{p}_{\mathrm{MA}}$ such that, for every $p^{\star}\in\{p_{1},\dots,p_{M}\}$, with probability at least $1-\delta$ over $X_{1},\dots,X_{n}$ i.i.d. from $p^{\star}$ one has
    $\operatorname{KL}({p^{\star}},{\widehat{p}_{\mathrm{MA}}})\leqslant\psi_{\mathrm{MA}}(M,n,\delta);$ (1)
  2. 2.
    there exists an estimator $\widehat{p}_{\mathrm{CA}}$ such that, for every $p^{\star}\in\mathop{\mathrm{conv}}\{p_{1},\dots,p_{M}\}$, with probability at least $1-\delta$ over $X_{1},\dots,X_{n}$ i.i.d. from $p^{\star}$ one has
    $\operatorname{KL}({p^{\star}},{\widehat{p}_{\mathrm{CA}}})\leqslant\psi_{\mathrm{CA}}(M,n,\delta).$ (2)
The fact that we make no assumption on the densities [^1] Given probability distributions $P_{1},\dots,P_{M}$ on $\mathcal{X}$, the existence of a dominating measure $\mu$ such that each $P_{j}$ admits a density $p_{j}$ with respect to $\mu$ is not restrictive, as one may simply take $\mu=P_{1}+\dots+P_{M}$. $p_{1},\dots,p_{M}$ induces some difficulties even for Model Aggregation. Indeed, in this case it is straightforward to show (see Fact 1 in Appendix B) that no selection method $\widehat{p}$ returning an element of the dictionary $\{p_{1},\dots,p_{M}\}$ achieves any meaningful distribution-free guarantee. In particular, this rules out arguably the simplest method, namely the maximum likelihood estimator (MLE). For this reason, the procedures we will consider will return mixtures of distributions from the dictionary, even for Model Aggregation.
Aside from the MLE, the most common method for Model Aggregation is Bayesian model averaging, which returns a mixture of the densities $p_{1},\dots,p_{M}$ weighted by their Bayesian posterior weights (say, starting from a uniform prior distribution over $p_{1},\dots,p_{M}$). In Appendix B, we show that this standard method is also suboptimal in the high-probability regime. Specifically, we obtain matching upper and lower bound on the tail performance of Bayes model averaging in the case where $M=2$, which show that this method achieves nontrivial yet suboptimal guarantees.
We note that there is a rich literature on Kullback-Leibler density estimation, both in the Model aggregation Barron (1987); Yang and Barron (1999); Yang (2000); Catoni (1997); Juditsky et al. (2008); Audibert (2007, 2009); Rigollet (2012); Butucea et al. (2017); Baraud (2021); van der Hoeven et al. (2023) and Convex aggregation Li and Barron (1999); Li (1999); Rakhlin et al. (2005); Dalalyan and Sebbar (2018) settings. However, our point of view differs from that of the aggregation literature. Indeed, a key emphasis in these works is the effect of model misspecification, where $p^{\star}$ does not belong to $\{p_{1},\dots,p_{M}\}$; while—with two exceptions discussed below—the ratios of densities $p_{1},\dots,p_{M}$ are assumed to be bounded, with the guarantees diverging when the bound on density ratios diverges. In contrast, we consider arbitrary densities $p_{1},\dots,p_{M}$, for which optimal high-probability guarantees are unknown even in the well-specified case (where $p^{\star}$ satisfies either of the two assumptions above). For these reasons, we focus in this work on optimal guarantees in the well-specified case, although we provide in Section 7 extensions to the misspecified case with additional logarithmic factors.
An important special case of mixture density estimation in which the boundedness condition fails to hold is that of estimation of discrete distributions, where one aims to estimate an arbitrary distribution on a finite set (say, $\{1,\dots,M\}$). This can be framed as mixture density estimation, where $\mathcal{X}=\{1,\dots,M\}$, $\mu$ is the counting measure (i.e., $\mu(A)$ is the cardinality of $A$ for $A\subset\mathcal{X}$), and for $1\leqslant j\leqslant M$ the density $p_{j}:\mathcal{X}\to\mathbb{R}^{+}$ is defined by $p_{j}(k)=1$ if $k=j$ and $p_{j}(k)=0$ if $k\neq j$. (In other words, any distribution on $\mathcal{X}$ is a mixture of Dirac masses.) In this case, it is relatively straightforward to show, and has been known for a long time Catoni (1997); Forster and Warmuth (2002); Braess and Sauer (2004); Mourtada and Gaïffas (2022), that optimal in-expectation guarantees are of order $M/n$. In contrast, the study of high-probability bounds is markedly more delicate and has received significant interest in recent years Bhattacharyya et al. (2021); Han et al. (2023); Canonne et al. (2023); van der Hoeven et al. (2023); Mourtada (2025); van der Hoeven et al. (2025), leading finally to optimal guarantees in Mourtada (2025). However, the techniques used in these works are largely specific to the discrete setting, and it is unclear how to extend them to the general mixture setting.
We now discuss existing guarantees in Model Aggregation and Convex Aggregation that apply to arbitrary densities $p_{1},\dots,p_{M}$—both of which only hold in expectation.
For Model Aggregation, optimal guarantees in expectation have been obtained in seminal work of Yang and Barron Yang and Barron (1999); Yang (2000) and independently by Catoni Catoni (1997, 2004). The corresponding estimator $\widetilde{p}$ (the progressive mixture rule) traces back to Barron (1987) and consists in a version of Bayes model averaging with additional averaging over sample sizes. For $M$ densities, it achieves the following guarantee:
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widetilde{p})\big]\;\leqslant\;\frac{\log M}{n+1}.$ (3)
While this conclusively addresses the important question of optimal in-expectation guarantees, it leaves the more delicate question of optimal high-probability guarantees unanswered. Specifically, the upper bound (3) implies that $\psi_{\mathrm{MA}}(M,n,\delta)\leqslant\log(M)/(n\delta)$, but this bound is highly suboptimal in the high-confidence regime of small $\delta$. In fact, as we discuss in Appendix B, the progressive mixture rule achieves suboptimal high-probability guarantees, hence another estimator $\widehat{p}$ is required.
For Convex Aggregation, optimal guarantees are unknown even in expectation. The best known general upper bound is achieved by a continuous version of the progressive mixture rule $\widetilde{p}$, applied to the probability simplex with uniform prior distribution. Combining regret bounds for the Bayes mixture method Hazan et al. (2007) (related log-loss regret guarantees for the simplex go back to Krichevsky–Trofimov Krichevsky and Trofimov (1981) and to the universal portfolio analysis of Cover Cover (1991), see also Jézéquel et al. (2025) for recent progress on this problem) with an online-to-batch conversion gives the following guarantee:
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widetilde{p})\big]\leqslant\frac{M\big(1+\log(1+n/M)\big)}{n+1}.$ (4)
This bound is known to be suboptimal at least in the canonical special case of discrete distribution estimation discussed above, due to the presence of the $\log(1+n/M)$ factor. In addition, as in the Model Aggregation case, it does not address the question of optimal high-probability guarantees.
1.1 Our contributions
Our main statistical contributions are optimal high-probability guarantees for both Model Aggregation and Convex Aggregation (as well as optimal in-expectation rates for Convex Aggregation), for arbitrary dictionary of densities. In what follows $a\lesssim b$ means that there exists an absolute constant $c>0$ such that $a\leqslant cb$.
  1. 1.
    Theorem 1: for any $M,n\geqslant 2$ and $\delta\in(e^{-n},e^{-1})$, the optimal high-probability rate (1) of Model Aggregation scales as
    $\psi_{\mathrm{MA}}(M,n,\delta)\asymp\frac{\log M\log(1/\delta)}{n}.$
  2. 2.
    Theorem 6: for any $n\geqslant M\geqslant 2$ and $\delta\in(e^{-n},e^{-1})$, the optimal high-probability rate (2) of Convex Aggregation scales as
    $\psi_{\mathrm{CA}}(M,n,\delta)\asymp\frac{M+\log M\log(1/\delta)}{n}.$
Specifically, the lower bounds above have been established in the special case of discrete distributions Mourtada (2025); van der Hoeven et al. (2025); our contribution is to establish matching upper bounds for arbitrary dictionaries.
A standard approach to these problems is to use metric entropy estimates: one derives statistical guarantees by analyzing coverings of the relevant class of distributions. As mentioned above, existing implementations of this strategy have not yielded sharp guarantees in our setting. In this work, we develop tools that make a metric-entropy approach sharp. There are two main obstacles: (i) we do not have sharp covering numbers for mixture models in KL divergence, and (ii) even with sharp covering numbers, the usual reductions would still not lead to sharp statistical bounds.
Our approach relies on three ingredients. First, for both problems we construct a preliminary estimator that is close to $p^{\star}$ in Hellinger distance. This metric is more amenable to entropy methods, but for the mixture case we need stronger local covering bounds than were previously available (to the best of our knowledge, this remained open; see Baraud and Birgé (2018) for the state of the art). We establish the required entropy estimate (Theorem 2) in Section 3. Second, we construct an $\varepsilon$-cover in KL divergence within the resulting local Hellinger ball. Third, we treat this KL cover as a finite set of candidates and feed it into the recent estimator of van der Hoeven et al. (2023) under bounded density ratios, which returns a distribution whose KL risk is nearly as good as the best candidate in the set. We briefly gloss over some additional technicalities. In particular, we mix candidates in a specific way to satisfy the conditions required by the estimator. For completeness, we provide a self-contained (less general) version of this estimator in Appendix˜A.
The second ingredient, namely constructing a KL cover of a local Hellinger ball, is a critical step of the proof. For reasons explained later, it reduces to the following more general geometric problem. Given a convex compact subset $K\subset\mathbb{R}_{+}^{d}$, we would like to find a finite subset $A\subset K$ with controlled cardinality satisfying the following (one-sided) multiplicative approximation guarantee: for every $\theta=(\theta_{j})_{1\leqslant j\leqslant d}\in K$, there exists an element $\phi=(\phi_{j})_{1\leqslant j\leqslant d}\in A$ such that $\theta_{j}\leqslant c\,\phi_{j}$ for every $j=1,\dots,d$, where $c>1$ is an absolute constant. We call such a set $A$ a $c$-ratio cover of $K$ (Definition 1).
In Section 4, we show that convex sets admit small ratio covers. In particular, we establish the following optimal cardinality estimate for ratio covers of compact convex sets, which is one of the main contributions of this work. We expect this covering result to be of independent interest, beyond its application to estimation of mixture distributions.
  1. 3.
    We show that there exist universal constants $C_{1},C_{2}>1$ such that the following holds: for every $d\geqslant 1$, every convex compact subset of $\mathbb{R}_{+}^{d}$ admits a $C_{1}$-ratio cover with at most $C_{2}^{d}$ elements (Theorem˜3).
With these ingredients in hand, we prove our high-probability guarantees for model aggregation (Section 2; using the first and third ingredients) and convex aggregation (Section 5; using all three ingredients).
1.2 Further applications of ratio covers
Theorem 3 also appears useful beyond our two main estimation results. We outline a few further applications here, and develop them in detail in Section 4.2 and Section 6.
Approximate Pareto curves.
Let $K\subset\mathbb{R}_{+}^{d}$ represent the set of attainable values in a multi-objective optimization problem. For instance, a point $x=(x_{1},\dots,x_{d})\in K$ could represent producing $x_{1}$ units of product 1, $x_{2}$ units of product 2, and so on. In general, the Pareto-optimal set may be infinite. The influential work of Papadimitriou and Yannakakis Papadimitriou and Yannakakis (2000) shows that there exists a small subset of $K$ that approximately dominates all points in $K$. In our terminology, a set that $(1+\varepsilon)$-dominates $K$ is exactly a $(1+\varepsilon)$-ratio cover. Their bounds for general $K$ depend on $d,\varepsilon$ and on the range parameter $R=\max_{i\in[d]}\frac{\max_{x\in K}x_{i}}{\min_{y\in K}y_{i}}$, where here and in what follows $d]=\{1,\ldots,d\}$. In Section 4.2, we use our ratio cover machinery (under convexity of $K$) to remove the dependence on $R$. We expect this to be useful in settings where $K$ is convex, for example when $K$ is the feasible region of a linear program.
Expected error guarantees via sharpening Yang–Barron.
As remarked earlier, Yang and Barron provide a general recipe for converting KL covering numbers into in-expectation error bounds. While very general, their method often introduces an unnecessary logarithmic factor. In Section 6.1, we sharpen the Yang–Barron guarantee in settings where the class is convex and amenable to our ratio-cover approach. The plan is: (i) obtain an estimator with squared Hellinger error $\varepsilon$ (often via entropy methods), (ii) use ratio covers to build a KL cover of the local Hellinger ball, and (iii) run the Yang–Barron estimator using this local KL cover. The key improvement is that we only need a KL cover locally, and the ratio-cover structure remains helpful even when the initial Hellinger estimator occasionally lands outside the target $\varepsilon$-ball (thanks to Lemma˜5). This yields our in-expectation guarantee for convex aggregation, scaling as M/n{M}/{n} up to constants.
KL covering numbers.
The recent PhD thesis “Divergence Coverings” of Tang Tang (2022) develops techniques for constructing KL covers and highlights the problem of bounding the KL covering number of the $d$-dimensional simplex. In Section 6.2, we review this line of work and show how ratio covers sharpen the KL covering number for the simplex. We emphasize that this section is mainly an illustration of how our techniques can sharpen recent results: in fact, our approach extends immediately beyond the simplex to the convex hull of an arbitrary dictionary of $d$ distributions, and can be used to control corresponding local covering numbers.
1.3 Related literature
The literature on aggregation under logarithmic loss is vast; we only highlight lines of work that are closest to our setting.
Sequential prediction, logarithmic loss, and universal coding.
Sequential prediction with logarithmic loss lies at the intersection of online learning, universal coding, and Bayesian mixture methods. In universal coding, predictors such as the Krichevsky–Trofimov estimator yield sharp nonasymptotic redundancy guarantees under log-loss Krichevsky and Trofimov (1981). In online learning, log-loss is mixable in the sense of Vovk Vovk (1990), which leads to logarithmic regret for exponential-weights predictors on the simplex; see, e.g., Vovk (1990, 1998); Cesa-Bianchi and Lugosi (2006); Hazan et al. (2007). Universal portfolio selection Cover (1991) provides another prominent instance: cumulative log-wealth is formally equivalent to log-loss prediction over the simplex. Parts of our analysis are closest to recent work converting sequential guarantees into high-probability statistical risk bounds for exp-concave losses via online-to-batch arguments van der Hoeven et al. (2023).
Statistical aggregation and progressive mixtures.
In model aggregation for density estimation, progressive mixture rules were introduced and analyzed in Yang (2000), with a closely related information-theoretic perspective in Yang and Barron (1999). These developments connect to earlier work on Bayes rules and information consistency Barron (1987) and to mixture-based approaches to model aggregation Catoni (1997), as well as to online viewpoints on aggregating strategies DeSantis et al. (1988); Vovk (1990). Mirror averaging, closely related to progressive mixtures, yields oracle inequalities in several aggregation setups (primarily for quadratic loss) Juditsky et al. (2008). The broader aggregation framework and minimax formulations of optimal rates were developed in Nemirovski (2000); Tsybakov (2003). For deviation bounds, progressive mixtures can be suboptimal in model aggregation settings (shown for square loss in Audibert (2007)), motivating alternative constructions not directly driven by regret; in particular, deviation-optimal aggregation for bounded regression losses has been obtained via $Q$-aggregation Dai et al. (2012); Lecué and Rigollet (2014); Bellec (2017); Mourtada et al. (2023).
Likelihood methods, entropy, and geometry.
A central tool in nonparametric density estimation is the control of likelihood-based procedures through metric entropy and likelihood-ratio inequalities, as in sieve MLE analyses Wong and Shen (1995) and model-selection/concentration methods Massart (2007). Estimator selection under Hellinger-type risks and test-based constructions has been developed in Baraud (2011) and refined in Baraud et al. (2017); Baraud and Birgé (2018). Closer to our focus, local (bracketing) geometric properties of finite mixture models are studied in Gassiat and Van Handel (2014) for finite location mixtures of a single base density under smoothness assumptions. In the special case of mixture density estimation over a finite dictionary, optimal KL aggregation rates for the MLE under boundedness assumptions on the dictionary are investigated in Li and Barron (1999); Li (1999); Rakhlin et al. (2005); Rigollet (2012); Butucea et al. (2017); Dalalyan and Sebbar (2018). By contrast, our results apply to arbitrary dictionaries of densities without imposing additional structural conditions.
1.4 Notation and preliminary results
Let $[k]$ denote $\{1,2,\ldots,k\}$. $\widehat{L}$ is the normalized negative log-likelihood, namely
$\widehat{L}(p)=-\frac{1}{n}\sum\limits_{i=1}^{n}\log p(X_{i}).$ (5)
Squared Hellinger divergence is denoted as
$\operatorname{H}^{2}(p,q)=\frac{1}{2}\int_{\mathcal{X}}\big(\sqrt{p(x)}-\sqrt{q(x)}\big)^{2}\mathrm{d}\mu(x).$ (6)
We will use the following relation between KL divergence and the Hellinger distance.
Lemma 1.
Assume that $p,q$ are two densities on $\mathcal{X}$ with respect to $\mu$. It holds that
$\operatorname{KL}(p,q)\leqslant\frac{2}{(\sqrt{e}-1)^{2}}\,\operatorname{H}^{2}(p,q)\,\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\right\}.$
For the sake of completeness we present the proof of this result in Appendix˜D. We note that similar identities appear in (Birgé and Massart, 1998, Lemma 5), (Birgé, 1983, Lemma 4.4) with the same constants appearing in (Tang, 2022, Fact 4), but for a definition of the Hellinger distance without $1/2$ in front. See also Birgé and Massart (1993).
Our next preliminary result is a corollary of Theorem 1 in van der Hoeven et al. (2023), which is stated there for more general exponentially concave losses. We present this result in the form required for our future analysis.
Lemma 2 (Model aggregation with bounded density ratios).
Assume $X_{1},\ldots,X_{n}$ are i.i.d. random variables with arbitrary and unknown density $p^{\star}$ on $(\mathcal{X},\mu)$. Let $p_{1},\ldots,p_{M}$ be densities such that for some $m>0$,
$\sup_{x\in\mathcal{X}}\max_{i,j\in[M]}\left|\log\left(\frac{p_{i}(x)}{p_{j}(x)}\right)\right|\leqslant m.$
There is an estimator $\widehat{p}$ such that for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\min_{j\in[M]}\operatorname{KL}({p^{\star}},{p_{j}})+\frac{2\log\left(M\right)+\frac{25}{4}\left(e-2\right)\left(m+1\right)\log\left(\frac{1}{\delta}\right)}{n}.$ (7)
Exact construction of the estimator as well as the self-contained proof of Lemma˜2 are presented in Appendix˜A. As a short comment, the estimator $\widehat{p}$ is a specific variance-adjusted version of the standard progressive mixture rule. Note that Lemma˜2 does not solve the model aggregation problem introduced above as it requires that the density ratios are bounded, an assumption we aim to avoid in this work.
1.5 Structure of the paper
In Section˜2 we study model aggregation over a finite family of densities and prove Theorem˜1. In Section˜3, we establish a sharp local covering bound in Hellinger distance for mixtures (Theorem˜2). Section˜4 is devoted to ratio covers, including the proof of Theorem˜3 and additional results on approximate Pareto curves. These tools are combined in Section˜5 to obtain the high-probability bound for convex aggregation (Theorem˜6). Further applications of ratio covers are presented in Section˜6: expected KL guarantees via a refined Yang–Barron argument (Section˜6.1) and KL covering numbers (Section˜6.2). We conclude with the misspecified setting in Section˜7.
The appendices contain a proof of the high-probability online-to-batch conversion (Appendix˜A), as well as lower bounds illustrating the suboptimality of classical estimators (Appendix˜B). In particular, we obtain matching upper and lower bounds on the tail behavior of Bayesian model averaging in the case of two densities, which falls short of the ideal deviation bound. Finally, we include a detailed presentation of the Birgé–Le Cam tournament (Appendix˜C).
2 Model aggregation
In this section, we provide a complete analysis of model aggregation over $M$ individual densities, where no assumptions are made on the individual densities. Despite its novelty, this result is technically less involved than the mixture case and can be seen as a preliminary building block introducing ideas that will also appear in this more complex setting. In particular, even at the preliminary stage, we do not need to use the Birgé-Le Cam tournaments, but rather exploit simple properties of the MLE in this case.
Theorem 1.
Let $X_{1},\ldots,X_{n}$ be an i.i.d. sample from some unknown density $p^{\star}$ that belongs to a known class of arbitrary densities $\mathcal{P}=\{p_{1},\ldots,p_{M}\}$ defined on the space $\mathcal{X}$ with respect to the measure $\mu$. There is a density estimator $\widehat{p}$ such that, with probability at least $1-\delta$,
$\operatorname{KL}\left(p^{\star},\widehat{p}\right)\leqslant\min\left\{\frac{75\log(2M)\log(2/\delta)}{n},\log(M)\right\}.$
This bound is optimal, in the sense that there are $M$ distributions such that no estimator can perform better up to universal constant factors, as shown in (Mourtada, 2025, Lemma 1) and (van der Hoeven et al., 2025, Theorem 11).
For the sake of presentation we assume that we are given a sample of size $2n$. In general, provided that the sample size is at least $2$ we can split the sample into two equally sized independent parts, while dropping at most one element of the original sample. This will only lead to a larger absolute constant in the final statement. We describe the estimator $\widehat{p}$ as follows:
  • Use the first $n$ observations (defining $\widehat{L}$) to construct a data-dependent set of almost empirical risk minimizers
    $\widehat{\mathcal{Q}}_{1}=\left\{p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}.$
  • Construct a (random) density class
    $\widehat{\mathcal{Q}}_{2}=\bigg\{\frac{1}{2}p+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q:p\in\widehat{\mathcal{Q}}_{1}\bigg\}.$ (8)
  • Using the second half of the sample, run the density estimator of Lemma˜2 over $\widehat{\mathcal{Q}}_{2}$ to output the final density $\widehat{p}$. Finally, if $75\log(2M)\log(2/\delta)\geqslant 2n\log(M)$ output instead $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$.
Let us quickly explain why the procedure works. We first show that the construction of the set $\widehat{\mathcal{Q}}_{1}$ of almost empirical likelihood maximizers provides a set of small Hellinger radius that contains $p^{\star}$. However, closeness in Hellinger radius does not imply closeness in the Kullback–Leibler divergence. This is where the smoothed class $\widehat{\mathcal{Q}}_{2}$ helps: the construction does not inflate the Hellinger distance, but it allows one to control the density ratio. However, as an inevitable price of smoothing $p^{\star}$ does not have to belong to $\widehat{\mathcal{Q}}_{2}$. This is where now the regret-based techniques of Lemma 2 help, since they provide both the optimal tail bounds and allow to bypass the problem $p^{\star}\notin\widehat{\mathcal{Q}}_{2}$. Remarkably, the final estimator is optimal up to multiplicative constant factors.
The proof of Theorem 1 requires a preliminary lemma. The first result is standard and similar derivations can be found in e.g. Wong and Shen (1995).
Lemma 3.
Let $\mathcal{P}=\{p_{1},\ldots,p_{M}\}$ be a finite set of densities with respect to a measure $\mu$ on $\mathcal{X}$. Fix $p^{\star}\in\mathcal{P}$ and let $X_{1},\ldots,X_{n}$ be i.i.d. distributed according to $p^{\star}$. For $\delta\in(0,1)$, set
$\widehat{\mathcal{Q}}_{1}=\bigg\{p\in\mathcal{P}:\ \widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(2M/\delta)}{n}\bigg\}.$
Then, with probability at least $1-\delta$, one has $p^{\star}\in\widehat{\mathcal{Q}}_{1}$ and
$\max_{p\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p^{\star},p)\leqslant\frac{2\log(2M/\delta)}{n}.$
Proof.
Fix $t>0$. For any $p\in\mathcal{P}$, by Markov’s inequality,
$\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(-nt/2)\right)$
$\leqslant\exp(nt/2)\,{\mathbf{E}}\left[\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\right]$
$=\exp(nt/2)\,\big(1-\operatorname{H}^{2}(p^{\star},p)\big)^{n}\leqslant\exp\Big(\tfrac{nt}{2}-n\operatorname{H}^{2}(p^{\star},p)\Big),$
where we used $1-u\leqslant\exp(-u)$. Let $\mathcal{P}_{t}=\{p\in\mathcal{P}:\operatorname{H}^{2}(p^{\star},p)\geqslant t\}$. By a union bound,
$\mathbf{P}\left(\exists p\in\mathcal{P}_{t}:\ \widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t\right)\leqslant\sum_{p\in\mathcal{P}_{t}}\exp(-nt/2)\leqslant M\,\exp(-nt/2).$
Note that for any $p\in\mathcal{P}$, $\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t$ implies $\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t$, since $\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\widehat{L}(p^{\star})$. Therefore,
$\mathbf{P}\left(\exists p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t\ \text{and}\ \operatorname{H}^{2}(p^{\star},p)\geqslant t\right)\leqslant M\exp(-nt/2).$
For the second part, apply the same argument with t-t in place of $t$. For any $p\in\mathcal{P}$, it holds that
$\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant-t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(nt/2)\right)$
$\leqslant\exp(-nt/2)\,\Big(\int\sqrt{p(x)p^{\star}(x)}\,d\mu(x)\Big)^{n}\leqslant\exp(-nt/2).$
By a union bound,
$\mathbf{P}\left(\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\widehat{L}(p^{\star})-t\right)\leqslant\sum_{p\in\mathcal{P}}\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant-t\right)\leqslant M\exp(-nt/2).$
Equivalently, we have
$\mathbf{P}\left(\widehat{L}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}(q)\geqslant t\right)\leqslant M\exp(-nt/2).$
Choose $t={2\log(2M/\delta)}/{n}$. Then each of the two bad events above has probability at most $\delta/2$, so by a union bound they fail simultaneously with probability at most $\delta$. Hence, with probability at least $1-\delta$, every $p\in\widehat{\mathcal{Q}}_{1}$ satisfies $\operatorname{H}^{2}(p^{\star},p)\leqslant t$, and also $\widehat{L}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t$ hence $p^{\star}\in\widehat{\mathcal{Q}}_{1}$. This yields the claim. ∎
Proof of Theorem 1.
Fix $\delta\in(0,1)$. By Lemma 3, with probability at least $1-\delta/2$, the set
$\widehat{\mathcal{Q}}_{1}=\left\{p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}$
contains $p^{\star}$ and has the Hellinger diameter at most $2\sqrt{{2\log(4M/\delta)}/{n}}$. From now on we work on this event and focus on the second half of the sample. Using the second half of the sample we output the density $\widehat{p}$ obtained using the procedure of Lemma˜2 with respect to $\widehat{\mathcal{Q}}_{2}$. Let
$\widetilde{p}=\frac{1}{2}p^{\star}+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q.$
[…]
$\sup_{p,q\in\widehat{\mathcal{Q}}_{2}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+|\widehat{\mathcal{Q}}_{1}|\bigr)\leqslant\log(2M).$
Therefore, by Lemma˜2 we have, with probability at least $1-\delta/2$,
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\operatorname{KL}({p^{\star}},{\widetilde{p}})+\frac{2\log(M)+\frac{25}{4}\left(e-2\right)\left(\log(2M)+1\right)\log\left({2}/{\delta}\right)}{n}.$ (9)
Finally, we have by Lemma 1 and the convexity of the Hellinger distance squared, observing that the ratio of $p^{\star}$ and $\frac{1}{2}p^{\star}+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q$ is bounded by $2$,
[…]
$\leqslant\frac{1}{(\sqrt{e}-1)^{2}}\operatorname{H}^{2}\left(p^{\star},\frac{1}{|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q\right)\leqslant\frac{2\log(4M/\delta)}{(\sqrt{e}-1)^{2}n}.$
Plugging this into (9) and using the union bound, we have with probability at least $1-\delta$,
[…]
$\leqslant\frac{25\log(2M)\log(2/\delta)}{n}.$
Note that we used a sample of size $2n$; for any original sample size at least $2$, splitting into two equal parts (dropping at most one observation) increases the absolute constant by at most a factor of $3$ (the worst case is a total sample size of $3$). Finally, if $75\log(2M)\log(2/\delta)\geqslant 2n\log(M)$, we output $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$, which guarantees $\operatorname{KL}(p^{\star},\widehat{p})\leqslant\log(M)$. The claim follows. ∎
3 Local Hellinger entropy for mixture models
In this section we study the local Hellinger geometry of convex hulls of densities. This result already suffices to establish the optimal rate for estimation of mixture distributions under Hellinger loss. By contrast, the best previously known statistical rates incur additional logarithmic factors due to reliance on VC-type covering arguments; see, for instance, (Baraud and Birgé, 2018, Proposition 7 and Corollary 3). Our estimation guarantee in Hellinger distance is given by the bound (13) below.
We first introduce notation. Let $M\geqslant 2$ and, as above, let $\mathcal{P}=\{p_{1},\dots,p_{M}\}$ be a set of densities on $\mathcal{X}$ with respect to $\mu$. Given a set of densities $\mathcal{P}$, define its local Hellinger entropy by
$N_{\mathsf{loc}}(\mathcal{P},\varepsilon)=\sup\limits_{p^{\star}\in\mathcal{P}}\sup_{\eta\geqslant\varepsilon}\ N_{\operatorname{H}}\big(B_{\operatorname{H}}(p^{\star},\eta)\cap\mathcal{P},\eta/2\big),$ (10)
where $B_{\operatorname{H}}(p,r)=\{q\text{ is a density on }(\mathcal{X},\mu):\operatorname{H}(p,q)\leqslant r\}$ and $N_{\operatorname{H}}(\mathcal{F},\varepsilon)$ is the covering number of $\mathcal{F}$ in the Hellinger metric $\operatorname{H}$ at scale $\varepsilon$. For any $\theta=(\theta_{1},\dots,\theta_{M})\in\Delta_{M-1}$, where in what follows
$\Delta_{M-1}=\bigg\{\theta\in\mathbb{R}^{M}_{+}:\sum_{j=1}^{M}\theta_{j}=1\bigg\},$
we define $p_{\theta}=\sum_{j=1}^{M}\theta_{j}p_{j}$. It is convenient to define the pseudo-metric $d_{\operatorname{H}}$ on the simplex $\Delta_{M-1}$, which is the Hellinger distance between mixtures, namely for any $\theta,\theta^{\prime}\in\Delta_{M-1}$, one has $d_{\operatorname{H}}(\theta,\theta^{\prime})=\operatorname{H}(p_{\theta},p_{\theta^{\prime}})$. The following result provides a universal upper bound on local Hellinger covering numbers of the class of mixtures of $M$ densities.
Theorem 2.
For any finite dictionary of densities $\mathcal{P}=\{p_{1},\dots,p_{M}\}$ on $\mathcal{X}$ with respect to the measure $\mu$ and any $\varepsilon>0$, one has
$N_{\mathsf{loc}}\big(\mathop{\mathrm{conv}}(\mathcal{P}),\varepsilon\big)\leqslant 64^{M-1}.$
Proof.
Fix $p^{\star}\in\mathop{\mathrm{conv}}(\mathcal{P})$ and $\eta\geqslant\varepsilon$, and choose $\theta_{0}\in\Delta_{M-1}$ such that $p^{\star}=p_{\theta_{0}}$. Set
$\mathcal{B}=\{\theta\in\Delta_{M-1}:d_{\operatorname{H}}(\theta,\theta_{0})\leqslant\eta\}.$
Then
$B_{\operatorname{H}}(p^{\star},\eta)\cap\mathop{\mathrm{conv}}(\mathcal{P})=\{p_{\theta}:\theta\in\mathcal{B}\},$
and moreover, for $A\subseteq\Delta_{M-1}$ and $\delta>0$, let $N_{d_{\operatorname{H}}}(A,\delta)$ denote the covering number of $A$ in the pseudo-metric $d_{\operatorname{H}}$ at scale $\delta$. Then
$N_{\operatorname{H}}\big(B_{\operatorname{H}}(p^{\star},\eta)\cap\mathop{\mathrm{conv}}(\mathcal{P}),\eta/2\big)\leqslant N_{d_{\operatorname{H}}}(\mathcal{B},\eta/2).$ (11)
Since $d_{\operatorname{H}}\leqslant 1$, if $\eta>1$ then $\mathcal{B}=\Delta_{M-1}$ and by monotonicity in the radius $N_{d_{\operatorname{H}}}(\mathcal{B},\eta/2)\leqslant N_{d_{\operatorname{H}}}(\Delta_{M-1},1/2)$, so it suffices to prove the bound below for $0<\eta\leqslant 1$.
Assume $0<\eta\leqslant 1$. By definition, $\mathcal{B}=\{\theta\in\Delta_{M-1}:\operatorname{H}^{2}(p_{\theta_{0}},p_{\theta})\leqslant\eta^{2}\}$; the map $\theta\mapsto p_{\theta}$ is affine and $q\mapsto\operatorname{H}^{2}(p_{\theta_{0}},q)$ is continuous and convex, so $\mathcal{B}$ is convex and compact. Moreover, for any $\theta,\theta^{\prime}\in\Delta_{M-1}$ and any $t\in[0,1]$, convexity of $\operatorname{H}^{2}(p_{\theta},\cdot)$ gives
$\operatorname{H}^{2}\left(p_{\theta},p_{(1-t)\theta+t\theta^{\prime}}\right)=\operatorname{H}^{2}\left(p_{\theta},(1-t)p_{\theta}+tp_{\theta^{\prime}}\right)\leqslant t\operatorname{H}^{2}(p_{\theta},p_{\theta^{\prime}}),$
and therefore
$d_{\operatorname{H}}\big(\theta,(1-t)\theta+t\theta^{\prime}\big)\leqslant\sqrt{t}\,d_{\operatorname{H}}(\theta,\theta^{\prime}).$ (12)
Using $d_{\operatorname{H}}\leqslant 1$ and (12) with $\theta=\theta_{0}$ shows that $(1-\eta^{2})\theta_{0}+\eta^{2}\Delta_{M-1}\subseteq\mathcal{B}$, hence (with $\mathrm{Vol}_{M-1}$ the Lebesgue measure on the affine hull $\{\theta\in\mathbb{R}^{M}:\sum_{j=1}^{M}\theta_{j}=1\}$) one has
$\mathrm{Vol}_{M-1}(\mathcal{B})\geqslant\mathrm{Vol}_{M-1}\big((1-\eta^{2})\theta_{0}+\eta^{2}\Delta_{M-1}\big)=\eta^{2(M-1)}\mathrm{Vol}_{M-1}(\Delta_{M-1})>0.$
Let $\{\theta_{1},\dots,\theta_{N}\}\subset\mathcal{B}$ be a maximal $(\eta/2)$-separated set in $(\mathcal{B},d_{\operatorname{H}})$. By the standard property we have $N_{d_{\operatorname{H}}}(\mathcal{B},\eta/2)\leqslant N$. Fix $\lambda=1/64$ and define $A_{i}=(1-\lambda)\theta_{i}+\lambda\mathcal{B}$ for $i=1,\dots,N$. Since $\mathcal{B}$ is convex and $\theta_{i}\in\mathcal{B}$, we have $A_{i}\subseteq\mathcal{B}$. If $\theta\in A_{i}$ and $\theta^{\prime}\in A_{j}$ with $i\neq j$, write $\theta=(1-\lambda)\theta_{i}+\lambda\widetilde{\theta}$ and $\theta^{\prime}=(1-\lambda)\theta_{j}+\lambda\widetilde{\theta}^{\prime}$ for some $\widetilde{\theta},\widetilde{\theta}^{\prime}\in\mathcal{B}$; then (12) with $t=\lambda$ and the triangle inequality yield
$d_{\operatorname{H}}(\theta_{i},\theta)\leqslant\sqrt{\lambda}\,d_{\operatorname{H}}(\theta_{i},\widetilde{\theta})\leqslant\sqrt{\lambda}\big(d_{\operatorname{H}}(\theta_{i},\theta_{0})+d_{\operatorname{H}}(\theta_{0},\widetilde{\theta})\big)\leqslant 2\eta\sqrt{\lambda},$
and similarly $d_{\operatorname{H}}(\theta_{j},\theta^{\prime})\leqslant 2\eta\sqrt{\lambda}$. Therefore,
$d_{\operatorname{H}}(\theta,\theta^{\prime})\geqslant d_{\operatorname{H}}(\theta_{i},\theta_{j})-d_{\operatorname{H}}(\theta_{i},\theta)-d_{\operatorname{H}}(\theta_{j},\theta^{\prime})>\frac{\eta}{2}-4\eta\sqrt{\lambda}=0,$
since $\sqrt{\lambda}=1/8$; hence $A_{i}\cap A_{j}=\emptyset$. Observe that $\mathrm{Vol}_{M-1}(A_{i})=\lambda^{M-1}\mathrm{Vol}_{M-1}(\mathcal{B})$. As the sets $A_{1},\dots,A_{N}$ are pairwise disjoint and contained in $\mathcal{B}$,
$\mathrm{Vol}_{M-1}(\mathcal{B})\geqslant\sum_{i=1}^{N}\mathrm{Vol}_{M-1}(A_{i})=N\,\lambda^{M-1}\mathrm{Vol}_{M-1}(\mathcal{B}),$
and since $\mathrm{Vol}_{M-1}(\mathcal{B})>0$ we conclude $N\leqslant\lambda^{-(M-1)}=64^{M-1}$. Thus $N_{d_{\operatorname{H}}}(\mathcal{B},\eta/2)\leqslant 64^{M-1}$, and combining with (11) gives
$N_{\operatorname{H}}\big(B_{\operatorname{H}}(p^{\star},\eta)\cap\mathop{\mathrm{conv}}(\mathcal{P}),\eta/2\big)\leqslant 64^{M-1}\quad\text{for all }p^{\star}\in\mathop{\mathrm{conv}}(\mathcal{P}),\ \eta\geqslant\varepsilon.$
Taking the suprema in (10) yields $N_{\mathsf{loc}}(\mathop{\mathrm{conv}}(\mathcal{P}),\varepsilon)\leqslant 64^{M-1}$. ∎
4 Ratio covers of convex sets
After approximating the target distribution in Hellinger distance (as we did above, for instance in the special case of finite mixtures), a central question in this work is how to convert this control into a small set of candidate distributions such that at least one is close to $p^{\star}$ in KL divergence. We recall from Lemma 1 that the KL divergence $\operatorname{KL}({p},{q})$ can be bounded in terms of the Hellinger distance together with a factor depending on the maximal density ratio $\sup_{x\in\mathcal{X}}\frac{p(x)}{q(x)}$. In particular, if $p$ and $q$ lie in the same squared Hellinger ball of radius $\varepsilon$ and the density ratio is bounded by a constant, then $\operatorname{KL}({p},{q})\leqslant C\,\varepsilon$ for some constant $C>0$ depending only on that ratio bound. This motivates the general notion of a ratio cover, which we will later apply to Hellinger balls to obtain the KL covers needed in our applications.
Definition 1.
Let $K$ be a compact subset of $\mathbb{R}_{+}^{d}$ and $\alpha\geqslant 1$. We say that a subset $A\subset K$ is an $\alpha$-ratio cover of $K$ if for every $\theta=(\theta_{1},\dots,\theta_{d})\in K$, there exists an element $\phi=(\phi_{1},\dots,\phi_{d})\in A$ such that $\theta_{j}\leqslant\alpha\,\phi_{j}$ for every $j=1,\dots,d$.
Our goal is to show the existence of $C_{1}$-ratio covers of size $(C_{2})^{d}$ for convex sets $K$, with constants $C_{1},C_{2}\geqslant 1$. In our proof of Theorem 6 for mixture models, we will ultimately pay a term proportional to $\log(|A|)/n$ for a ratio cover of size $|A|$: hence even a ratio cover of size $|A|=(\log d)^{d}$ would yield a suboptimal upper bound featuring an additional $\log\log d$ factor, and similarly any super-constant ratio $\alpha$ would yield suboptimal results (both issues are also important for our other applications). With this motivation, we state one of our main results.
Theorem 3.
For every $d\geqslant 1$ and every convex and compact set $K\subset\mathbb{R}_{+}^{d}$, there exists a subset $A\subset K$ with at most $2^{8d}$ elements that is a $32$-ratio cover of $K$.
In the remainder of this section, we first briefly comment on the theorem statement, and then we discuss an idea of the proof of Theorem 3. In Subsection 4.1, we prove Theorem 3.
First, the constraint $A\subset K$ is essential: without it, there is a trivial “cover” obtained by choosing the single vector $\phi\in\mathbb{R}_{+}^{d}$ with coordinates $\phi_{j}=\max_{\theta\in K}\theta_{j}$, which need not belong to $K$ and is not useful for our applications. Second, the definition must be one-sided. Indeed, if one required the two-sided condition $\phi_{j}/\alpha\leqslant\theta_{j}\leqslant\alpha\phi_{j}$, then even in dimension $d=1$ the set $K=[1,L]$ with $L\geqslant 2$ would require at least $\log_{\alpha}L$ points. Third, convexity is required: for $d=2$, the nonconvex set
$K=\{(x,1/x):\ 1/L\leqslant x\leqslant L\}$
also requires at least $\log_{\alpha}L$ points for any fixed $\alpha$. (Of course, one could drop the convexity assumption on $K$ but then relax the requirement to $A\subset\mathop{\mathrm{conv}}(K)$.) Finally, even for convex $K$, an exponential dependence on $d$ is sometimes unavoidable for constant $\alpha$. A simple illustration is $\Delta_{1}=\{(t,1-t):t\in[0,1]\}\subset\mathbb{R}_{+}^{2}$: if $\alpha=1.9$, no single point $\phi=(a,1-a)\in\Delta_{1}$ can $\alpha$-ratio cover both extreme points $(1,0)$ and $(0,1)$. Hence any $1.9$-ratio cover of $\Delta_{1}$ must have size at least $2$. Consequently, if $d$ is even, a $1.9$-ratio cover for the product set $\prod_{i=1}^{d/2}\Delta_{1}\subset\mathbb{R}_{+}^{d}$ must have size at least $2^{d/2}$: each block $\Delta_{1}$ contributes an independent binary choice, so one needs a distinct cover point for each of the $2^{d/2}$ combinations.
We provide two different approaches to the proof of Theorem˜3. The first is somewhat longer, but fully constructive and deterministic, based on first principles and exploiting the convexity of the set. The second proof, presented in Section˜4.3, is shorter and relies on a reduction to $\varepsilon$-net arguments presented in Diakonikolas and Yannakakis (2010) together with the volumetric argument used in the proof of Theorem˜2.
Proof roadmap for Theorem˜3.
We outline the main ideas of the proof. We begin with two warm-ups that highlight the basic geometry behind ratio covers, and then explain how the full argument combines discretization, averaging, and induction to obtain a $C_{1}$-ratio cover of size $C_{2}^{d}$.
Warm-up 1: a $d$-ratio cover with a single point. For each coordinate $j\in[d]$, let $m_{j}=\max_{\theta\in K}\theta_{j}$ (the maximum exists since $K$ is compact), and pick $\overline{\theta}^{(j)}\in K$ such that $\overline{\theta}^{(j)}_{j}=m_{j}$. Define
$\phi^{(1)}=\frac{1}{d}\sum_{j=1}^{d}\overline{\theta}^{(j)}\in K,$
where $\phi^{(1)}\in K$ by convexity. Then for any $\theta\in K$ and any $i\in[d]$,
$\theta_{i}\leqslant m_{i}=\overline{\theta}^{(i)}_{i}\leqslant\sum_{j=1}^{d}\overline{\theta}^{(j)}_{i}=d\,\phi^{(1)}_{i}$.
Thus $\{\phi^{(1)}\}$ is a $d$-ratio cover of $K$.
Warm-up 2: a $4$-ratio cover with $|A|\leqslant(\lceil\log_{2}d\rceil+1)^{d}$. Warm-up 1 is too coarse since the ratio factor grows with $d$. To reduce the ratio, we discretize each coordinate of $\theta$ into dyadic bins relative to $\phi^{(1)}$. Fix $\theta\in K$. If $\theta_{i}\leqslant\phi^{(1)}_{i}$, then coordinate $i$ is already $1$-covered by $\phi^{(1)}$. Otherwise, $\theta_{i}>\phi^{(1)}_{i}$, and the bound from Warm-up 1 implies $\theta_{i}\leqslant d\,\phi^{(1)}_{i}$. (If $\phi^{(1)}_{i}=0$, then necessarily $m_{i}=0$ and hence $\theta_{i}=0$ for all $\theta\in K$, so only the first case can occur.) Therefore each coordinate $i$ must satisfy either
$\theta_{i}\leqslant\phi^{(1)}_{i},\quad\text{or}\quad\theta_{i}\in(2^{j}\phi^{(1)}_{i},\,2^{j+1}\phi^{(1)}_{i}]\ \text{for some }j\in\{0,\dots,\lceil\log_{2}d\rceil-1\}.$
This yields at most $(\lceil\log_{2}d\rceil+1)^{d}$ possible bin patterns across all coordinates. Let $A_{2}\subset K$ contain one representative $\phi^{(2)}$ for each bin pattern that occurs among points of $K$, and define
$A=\frac{\phi^{(1)}+A_{2}}{2}=\left\{\frac{\phi^{(1)}+\phi^{(2)}}{2}:\ \phi^{(2)}\in A_{2}\right\}.$
Given $\theta\in K$, pick $\phi^{(2)}\in A_{2}$ with the same bin pattern. If $\theta_{i}\leqslant\phi^{(1)}_{i}$, then
$\theta_{i}\leqslant\phi^{(1)}_{i}\leqslant\phi^{(1)}_{i}+\phi^{(2)}_{i}=2\frac{\phi^{(1)}_{i}+\phi^{(2)}_{i}}{2}.$
Otherwise, for some $j$ we have $\theta_{i}\in(2^{j}\phi^{(1)}_{i},2^{j+1}\phi^{(1)}_{i}$], and the matching bin pattern implies $\phi^{(2)}_{i}>2^{j}\phi^{(1)}_{i}$, hence
$\theta_{i}\leqslant 2^{j+1}\phi^{(1)}_{i}\leqslant 2\,\phi^{(2)}_{i}\leqslant 4\frac{\phi^{(1)}_{i}+\phi^{(2)}_{i}}{2}.$
Therefore $A$ is a $4$-ratio cover of $K$, and $|A|=|A_{2}|\leqslant(\lceil\log_{2}d\rceil+1)^{d}$. We note that the idea of discretizing by dyadic scales and selecting one representative per pattern also appears in Papadimitriou and Yannakakis (2000).
Idea of the full theorem. The warm-ups still fall short: we need a constant ratio $C_{1}$ and cover size $C_{2}^{d}$. After discretizing $K$ as above, the key additional observation is that averaging a finite set of discretized representatives typically provides slack on many coordinates: for a large fraction of discretized points, the average already constant-covers at least half of the coordinates. This enables the following high-level strategy:
  1. 1.
    Discretize points of $K$ into finitely many profiles (as in Warm-up 2), and pick one representative per profile.
  2. 2.
    Iteratively take averages of the remaining representatives; at each step, keep the average and discard all representatives that are already constant-covered by it on at least half of the coordinates.
  3. 3.
    For the uncovered coordinates (at most half), apply the induction hypothesis to a suitable lower-dimensional projection.
  4. 4.
    Combine the averaging-based candidates with the inductive covers so that every coordinate is covered with an overall constant ratio.
4.1 Constructive proof of Theorem 3
We proceed by induction on $d\geqslant 1$.
Base case $d=1$.
This case follows simply. Let $\theta^{(0)}=\mathop{\mathrm{arg}\,\mathrm{max}}_{\theta\in K}\theta$ (which exists by compactness of $K$). Then, $A=\{\theta^{(0)}\}\subset K$ is a $1$-ratio cover of $K$ with one element.
Induction.
Otherwise, let $d\geqslant 2$; we assume that the property holds for any dimension $d^{\prime}<d$, and proceed to show it in dimension $d$.
We will define below four finite subsets $A_{1},\dots,A_{4}\subset K$ such that
$A\subset(A_{1}+\dots+A_{4})/4=\{(\phi^{(1)}+\dots+\phi^{(4)})/4:\phi^{(1)}\in A_{1},\dots,\phi^{(4)}\in A_{4}\}$
is a $32$-ratio cover of $K$.
Definition of $A_{1}$. $A_{1}$ is constructed exactly as in the first warm-up. For each $j=1,\dots,d$, define $m_{j}=\sup_{\theta=(\theta_{1},\dots,\theta_{d})\in K}\theta_{j}$, which is finite and attained by compactness of $K$. Then, for each $j=1,\dots,d$, pick $\overline{\theta}^{(j)}=(\overline{\theta}_{1}^{(j)},\dots,\overline{\theta}_{d}^{(j)})\in K$ such that $\overline{\theta}_{j}^{(j)}=m_{j}$, and define $\phi^{(1)}=(\overline{\theta}^{(1)}+\dots+\overline{\theta}^{(d)})/d\in K$. We then simply take $A_{1}=\{\phi^{(1)}\}$.
For every $\theta=(\theta_{j})_{1\leqslant j\leqslant d}\in K$, define
$J_{1}(\theta)=\{1\leqslant j\leqslant d:\theta_{j}\leqslant m_{j}/d\}.$
Then, by definition of $\phi^{(1)}$, for each $j\in J_{1}(\theta)$ one has $\theta_{j}\leqslant m_{j}/d\leqslant\phi_{j}^{(1)}$, i.e., the $j$-th coordinate of $\theta$ is $1$-covered by that of $\phi^{(1)}$.
Definition of $A_{2}$. We now proceed to cover some of the remaining coordinates. For every subset $J\subseteq[d]$, define
$K_{1}(J)=\{\theta\in K:[d]\setminus J_{1}(\theta)=J\},$
the set of vectors in $K$ whose set of “remaining coordinates” is $J$. We will now define a subset $A_{2}(J)\subset K$ that covers at least half of the coordinates for every vector $\theta\in K_{1}(J)$.
If $J=\emptyset$ (meaning that there are no coordinates left to cover), then arbitrarily set $A_{2}(J)=\{\phi^{(1)}\}$; if $K_{1}(J)=\emptyset$ (meaning that there are no vectors to cover), then set $A_{2}(J)=\emptyset$. Now, let $J\neq\emptyset$ such that $K_{1}(J)\neq\emptyset$. By definition of $K_{1}(J)$, for every $\theta\in K_{1}(J)$ and $j\in J$, we know $m_{j}/d<\theta_{j}\leqslant m_{j}$.
We first discretize the vectors $\{(\theta_{j})_{j\in J}:\theta\in K_{1}(J)\}$ by powers of $2$ of their coordinates. Let us formalize this: we define the profile of a vector $\theta\in K_{1}(J)$ as the vector
$(k_{j})_{j\in J}\in\mathbb{N}^{J}\textrm{ such that }m_{j}2^{-k_{j}-1}<\theta_{j}\leqslant m_{j}2^{-k_{j}}\textrm{ for every }j\in J.$
Since $m_{j}/d<\theta_{j}\leqslant m_{j}$, one has $0\leqslant k_{j}\leqslant\left\lceil\log_{2}d\right\rceil-1\leqslant\log_{2}d$.
We say that a profile is admissible if an element of $K_{1}(J)$ has this profile. We then define a subset $S\subset K_{1}(J)$ containing exactly one representative of each admissible profile. Note that the cardinality of $S$ is at most $|S|\leqslant|\{0,\dots,\left\lceil\log_{2}d\right\rceil-1\}^{J}|\leqslant(\log_{2}d+1)^{d}$.
So far, we have produced a finite set of representatives that represent the (possibly infinite) remaining values of $\theta$. Since each $\theta$ has its coordinates approximated by its representative, our intuition is that producing a ratio-cover for this set of representatives will yield a ratio-cover for all $\theta$. In this sense, we have essentially reduced the task of ratio-covering a convex set $K$, to the task of providing a ratio cover for a set with bounded size. At first, this may not seem much easier; particularly because the set is still quite large. However, let us first consider a weaker goal, where we only ratio-cover half of the coordinates:
Definition 2 (Half-$c$-cover).
For $c\geqslant 1$, we say that a vector $\theta$ is half-$c$-covered by a vector $\mu$ if the inequality $\theta_{j}\leqslant c\mu_{j}$ holds for at least half of the indices $j\in J$.
Even producing a half-$c$-cover for our bounded-size set is not immediately obvious. However, we will argue that the average of the set $S$ is a half-$4$-cover for at least half of the elements in $S$:
Claim 1.
Given a non-empty set of indices $J\subseteq[d]$, let $S^{\prime}\subset\mathbb{R}_{+}^{d}$ be a finite set, and define $\mu^{\prime}=\frac{1}{|S^{\prime}|}\sum_{\theta^{\prime}\in S^{\prime}}\theta^{\prime}$. Then, at least half of the elements of $S^{\prime}$ are half-$4$-covered by $\mu^{\prime}$.
Proof of Claim 1.
Denote by $B\subset S^{\prime}$ the set of $\theta^{\prime}\in S^{\prime}$ that are not half-$4$-covered by $\mu^{\prime}$. For each $j\in J$, Markov’s inequality implies that $|\{\theta\in S^{\prime}:\theta_{j}>4\mu^{\prime}_{j}\}|\leqslant|S^{\prime}|/4$. Hence,
$\frac{|S^{\prime}|}{4}\geqslant\frac{1}{|J|}\sum_{j\in J}\sum_{\theta\in S^{\prime}}\bm{1}(\theta_{j}>4\mu^{\prime}_{j})=\sum_{\theta\in S^{\prime}}\frac{\lvert\{j\in J:\theta_{j}>4\mu_{j}^{\prime}\}\rvert}{|J|}\geqslant|B|\cdot\frac{|J|/2}{|J|}=\frac{|B|}{2}\,$,
so that $|B|\leqslant|S^{\prime}|/2$, which concludes the proof. ∎
Our plan is to invoke this claim repeatedly to produce a half-$c$-cover for all of $S$. Concretely, we recursively define elements $\mu^{(k)}\in K$ and subsets $S_{k}\subset S$ as follows. Let $S_{1}=S$. Then, for each $k\geqslant 1$:
  • Define $\mu^{(k)}=\frac{1}{|S_{k}|}\sum_{\theta\in S_{k}}\theta\in K$.
  • Let $S_{k+1}$ be the set of vectors in $S_{k}$ that are not half-$4$-covered by $\mu^{(k)}$.
  • If $S_{k+1}=\emptyset$, then stop; otherwise, turn to step $k+1$.
By Claim 1, one has $|S_{k+1}|\leqslant|S_{k}|/2$, hence the loop stops after
$r\leqslant\left\lfloor\log_{2}|S|\right\rfloor+1\leqslant d\log_{2}(\log_{2}d+1)+1\leqslant 2^{d}$
iterations. We define $A_{2}(J)=\{\mu^{(1)},\dots,\mu^{(r)}\}$, the set of partial means, such that $|A_{2}(J)|\leqslant 2^{d}$ by the above.
By construction of the means $\mu^{(k)}$ and sets $S_{k}$ above, each $\theta\in S$ is half-$4$-covered by an element of $A_{2}(J)$. In addition, by definition of $S$, each vector $\theta\in K_{1}(J)$ has all its coordinates $j\in J$ within a factor of $2$ of those of an element in $S$. Hence, $\theta$ is also half-$8$-covered by an element $\phi^{(2)}=\phi^{(2)}(\theta)\in A_{2}(J)$. We denote the coordinates handled by this step by
$J_{2}(\theta)=\{j\in J:\theta_{j}\leqslant 8\phi_{j}^{(2)}\}.$
Since this covers at least half of the coordinates for any $\theta$, we know $|J_{2}(\theta)|\geqslant|J|/2$. Thus, the number of remaining unhandled coordinates is at most
$|[d]\backslash(J_{1}(\theta)\cup J_{2}(\theta))|\leqslant d-|J_{1}(\theta)|-|[d]\backslash J_{1}(\theta)|/2\leqslant d/2.$
In summary, define $A_{2}=\bigcup_{J\subseteq[d]}A_{2}(J)$, such that $|A_{2}|\leqslant\sum_{J\subseteq[d]}2^{d}=2^{2d}$. Since $K=\bigcup_{J\subseteq[d]}K_{1}(J)$, it follows from the above that for every $\theta\in K$, there exists $\phi^{(2)}=\phi^{(2)}(\theta)\in A_{2}$ such that all coordinates $j\in J_{2}(\theta)$ of $\theta$ are $8$-covered by the corresponding ones of $\phi^{(2)}$.
Definition of $A_{3}$. For any $\theta\in K$, let $J^{\prime}(\theta)$ denote the coordinates not handled by $A_{1}$ and $A_{2}$:
$J^{\prime}(\theta)=[d]\backslash(J_{1}(\theta)\cup J_{2}(\theta)).$
As noted before, the number of unhandled coordinates is at most $d/2$. In what follows, we will handle the remaining coordinates $J^{\prime}(\theta)$ by induction, together with an argument to avoid degrading the ratio factor.
For every subset $J\subset[d]$ with $|J|\leqslant d/2$, we define $\pi_{J}(K)$ as the projection of $K$ to the coordinates $J$, meaning
$\pi_{J}(K)=\{(\theta_{j})_{j\in J}:(\theta_{j})_{1\leqslant j\leqslant d}\in K\}.$
The set $\pi_{J}(K)$ is a convex compact subset in dimension $d^{\prime}\leqslant d/2<d$; hence, by the induction hypothesis, it admits a subset with at most $2^{8d^{\prime}}\leqslant 2^{4d}$ elements that is a $32$-ratio cover.
We define the corresponding subset $A_{3}(J)\subset K$, with at most $2^{4d}$ elements, where $\pi_{J}(A_{3}(J))$ is a 32-ratio cover of $\pi_{J}(K)$. For $J=\emptyset$, we set $A_{3}(\emptyset)=\{\phi^{(1)}\}$. By the ratio cover property, for every $\theta\in K$, there exists $\phi^{(3)}=\phi^{(3)}(\theta)\in A_{3}(J^{\prime}(\theta))$ such that $\theta_{j}\leqslant 32\,\phi_{j}^{(3)}$ for every $j\in J^{\prime}(\theta)$. From here, we define $A_{3}=\bigcup_{J\,:\,|J|\leqslant d/2}A_{3}(J)$, with cardinality $|A_{3}|\leqslant\sum_{J\,:\,|J|\leqslant d/2}|A_{3}(J)|\leqslant 2^{d}\times 2^{4d}=2^{5d}$.
At this point, it is worth noting that achieving a $32$-covering from $\phi^{(3)}$ over the coordinates $j\in J^{\prime}(\theta)$ is not sufficient for our purposes, since the vector $\phi^{(3)}$ is to be averaged with previous vectors $\phi^{(1)},\phi^{(2)}$ to cover coordinates $j\in[d]\setminus J^{\prime}(\theta)$. This averaging operation will degrade the constant $32$ in the ratio factor, thus breaking the induction.
In order to address this issue, we distinguish between two types of coordinates. For $\theta\in K$, let $\phi^{(3)}=\phi^{(3)}(\theta)\in A_{3}(J^{\prime}(\theta))$ be as above. We define
$J_{3}(\theta)=\{j\in J^{\prime}(\theta):\theta_{j}\leqslant 8\,\phi_{j}^{(3)}\}\textrm{ and }J_{4}(\theta)=J^{\prime}(\theta)\setminus J_{3}(\theta)=[d]\setminus(J_{1}(\theta)\cup J_{2}(\theta)\cup J_{3}(\theta)).$
By definition, coordinates $j\in J_{3}(\theta)$ are $8$-covered, which provides the necessary slack to maintain a $32$-ratio covering after the averaging operation. It thus remains to handle the coordinates in $J_{4}(\theta)$.
Definition of $\phi^{(4)}$. We will not actually define a set $A_{4}$ as our earlier definition suggests, but instead will define a function $\phi^{(4)}(\theta)$ as before.
The final idea of the proof is the following: coordinates $j\in J_{4}(\theta)$ are $32$-covered but not $8$-covered by $\phi^{(3)}(\theta)$; meaning, the value of $\theta_{j}$ is known up to a factor of $4$, and can thus be separately encoded. Specifically, for any $\phi^{(3)}\in A_{3}$ and subset $J\subset[d]$, let
$C(\phi^{(3)},J)=\{\theta\in K:8\phi_{j}^{(3)}<\theta_{j}\leqslant 32\phi_{j}^{(3)},\ \textrm{for all}\ j\in J\}.$
Using this, let $R(\phi^{(3)},J)\in C(\phi^{(3)},J)$ be an arbitrary representative chosen for $C(\phi^{(3)},J)$ (as long as the set is nonempty). For any $\theta\in K$, let $\phi^{(3)}=\phi^{(3)}(\theta)$ and $J_{4}=J_{4}(\theta)$ as above. By definition of $J_{4}(\theta)$, we know $\theta\in C(\phi^{(3)},J_{4})$; hence, the latter set is nonempty, and we choose $\phi^{(4)}(\theta)=R(\phi^{(3)},J_{4})$. This ensures that $\theta_{j}\leqslant 32\phi_{j}^{(3)}\leqslant 4\phi_{j}^{(4)}$ for every $j\in J_{4}$.
Conclusion. We let
$A=\bigg\{\frac{\phi^{(1)}+\phi^{(2)}(\theta)+\phi^{(3)}(\theta)+\phi^{(4)}(\theta)}{4}:\theta\in K\bigg\}\,$,
which is contained in $K$ by convexity. In addition, for every $\theta\in K$, letting $\phi=(\phi^{(1)}+\phi^{(2)}(\theta)+\phi^{(3)}(\theta)+\phi^{(4)}(\theta))/4=(\phi_{j})_{1\leqslant j\leqslant d}$ and $J_{k}=J_{k}(\theta)$ for $k=1,\dots,4$, we have by the above that:
[…]
Since $d]=J_{1}\cup J_{2}\cup J_{3}\cup J_{4}$, we deduce that $A$ is a $32$-ratio cover of $K$.
It remains to control the cardinality of $A$. There are at most $|A_{2}|\leqslant 2^{2d}$ possible values of $\phi^{(2)}(\theta)\in A_{2}$, and at most $|A_{3}|\leqslant 2^{5d}$ possible values for $\phi^{(3)}(\theta)\in A_{3}$; in addition, since $\phi^{(4)}(\theta)=R(\phi^{(3)}(\theta),J)$ for some $J\subset[d]$, there are at most $2^{d}$ possible values of $\phi^{(4)}(\theta)$ given $\phi^{(3)}(\theta)$. Hence, there are at most $2^{2d}\times 2^{5d}\times 2^{d}\leqslant 2^{8d}$ possible values for $(\phi^{(2)}(\theta),\phi^{(3)}(\theta),\phi^{(4)}(\theta))$ for $\theta\in K$, and therefore $|A|\leqslant 2^{8d}$. ∎
4.2 Approximate Pareto curves
Consider a setting where a set $K\subset\mathbb{R}_{+}^{d}$ represents the attainable objective values in a multi-objective optimization problem. For instance, one may think of producing $d$ different products: a point $x=(x_{1},\dots,x_{d})\in K$ indicates that it is possible to simultaneously produce $x_{j}$ units of product $j$ for each $j\in[d]$. Such problems typically involve tradeoffs between objectives, and a common object of interest is the Pareto curve (or Pareto frontier), namely the set of points in $K$ that are not dominated coordinatewise by any other point in $K$. In general, the Pareto curve may be infinite. Papadimitriou and Yannakakis Papadimitriou and Yannakakis (2000) showed that one can nevertheless extract a finite $\varepsilon$-approximate Pareto curve, meaning a subset $A\subset K$ such that for every $\theta=(\theta_{1},\dots,\theta_{d})\in K$ there exists $\phi=(\phi_{1},\dots,\phi_{d})\in A$ with
$\theta_{j}\leqslant(1+\varepsilon)\,\phi_{j}\quad\text{for all }j\in[d].$
In our terminology, this is exactly a $(1+\varepsilon)$-ratio cover of $K$. Their result does not assume convexity of $K$, but it does assume that each coordinate ranges over a bounded interval.
Theorem 4 (Rephrased Theorem 1 of Papadimitriou and Yannakakis Papadimitriou and Yannakakis (2000)).
For every set $K\subset[\frac{1}{R},R]^{d}$, every $0<\varepsilon\leqslant 1$, and every $R\geqslant 1$, there exists a subset $A\subset K$ with
$|A|\leqslant\max\left\{1,\left(c\,\frac{\log R}{\varepsilon}\right)^{d-1}\right\}$
that is an $\varepsilon$-approximate Pareto curve (equivalently, a $(1+\varepsilon)$-ratio cover) for $K$, where $c>0$ is a universal constant.
Their short proof is essentially the same dyadic discretization idea that we used in our second warm-up. A key feature of Theorem˜4 is its dependence on the range parameter $R$. If one adds the assumption that $K$ is convex, then the averaging idea from our first warm-up yields a simple way to remove this dependence.
Corollary 1 (Corollary of Theorem˜4 under convexity).
For every convex and compact set $K\subset\mathbb{R}_{+}^{d}$ and every $0<\varepsilon\leqslant 1$, there exists a subset $A\subset K$ with
$|A|\leqslant\left(c\,\frac{\log d}{\varepsilon}\right)^{d-1}$
that is an $\varepsilon$-approximate Pareto curve for $K$, where $c>0$ is a universal constant.
By comparison, we now use our ratio cover theorem to obtain an $\varepsilon$-approximate Pareto curve whose size has no dependence on $R$ and whose base does not involve $d$. The convexity assumption is essential for this stronger statement (see the discussion following Theorem˜3), and it is natural in many applications, for example when $K$ is the feasible region of a linear program.
Corollary 2 (of Theorem˜3).
For every convex and compact set $K\subset\mathbb{R}_{+}^{d}$ and every $0<\varepsilon\leqslant 1$, there exists an $\varepsilon$-approximate Pareto curve $A\subset K$ with
$|A|\leqslant 2^{8d}\left(3+\frac{4\log_{2}(128/\varepsilon)}{\varepsilon}\right)^{d-1}.$
The proof of Corollary˜2 is deferred to Appendix˜D. A substantial literature following Papadimitriou and Yannakakis develops algorithms that construct $\varepsilon$-approximate Pareto curves of near-minimal size (see, e.g., Vassilvitskii and Yannakakis (2005); Diakonikolas and Yannakakis (2008, 2010); Bazgan et al. (2015); Daskalakis et al. (2016)). Our result shows that under convexity, there always exists an $\varepsilon$-approximate Pareto curve of smaller size, and therefore yields improved existential upper bounds for what such approximation algorithms can aim to achieve in this setting.
4.3 Proof of a version of Theorem˜3 based on the reduction to $\varepsilon$-nets
In this section, we provide an alternative proof of Theorem˜3 based on $\varepsilon$-net arguments in the spirit of Haussler and Welzl Haussler and Welzl (1986), introduced in the context of Pareto curves by Diakonikolas and Yannakakis Diakonikolas and Yannakakis (2010) for general sets $K$. We then combine this approach with convexity via a volumetric argument similar to the one used in the proof of Theorem˜2. We state the resulting bound as a separate theorem because it yields a different dependence on $\varepsilon$ in the language of Corollary˜2.
Theorem 5.
For every convex and compact set $K\subset\mathbb{R}_{+}^{d}$ with nonempty interior and every $\varepsilon>0$, there exists an $\varepsilon$-approximate Pareto curve $A\subset K$ with
$|A|\leqslant\left\lceil 8d\left(1+\frac{1}{\varepsilon}\right)^{d}\log\left(13\left(1+\frac{1}{\varepsilon}\right)^{d}\right)\right\rceil.$
Note that, in full generality, the dependence on $\varepsilon$ in Theorem˜5 differs from that in Corollary˜2. For instance, in the regime of fixed $d$ and sufficiently small $\varepsilon$, the bound of Corollary˜2 is stronger because it has exponent d1d-1 rather than $d$. Despite not having the same dependence on $\varepsilon$ as Corollary˜2, the result of Theorem˜5 implies a $32$-ratio cover of size at most
$\left\lceil 8d\left(\frac{32}{31}\right)^{d}\log\left(13\left(\frac{32}{31}\right)^{d}\right)\right\rceil\leqslant 2^{5d},$
for all $d\geqslant 1$. Since optimization of numerical constants is not our priority, we mostly use the simple integer form of Theorem˜3.
We also note that the construction in our first proof of Corollary˜2 is deterministic and constructive, whereas the proof of Theorem˜5 is nonconstructive and based on a probabilistic argument.
Proof.
The nonempty interior assumption implies $\mathrm{Vol}(K)>0$. For each $\theta\in K$, set
$R_{\theta}(\varepsilon)=\{\phi\in K:\theta_{j}\leqslant(1+\varepsilon)\phi_{j}\ \textrm{for all}\ j\in[d]\}.$
As observed in (Diakonikolas and Yannakakis, 2010, Section 4.1), $A$ is an $\varepsilon$-approximate Pareto curve if and only if, for all $\theta\in K$,
$R_{\theta}(\varepsilon)\cap A\neq\emptyset.$
Indeed, this means that for any $\theta\in K$ there exists $\phi\in A$ such that $\theta_{j}\leqslant(1+\varepsilon)\phi_{j}$ for all $j\in[d]$.
Now define the collection of sets $\mathcal{R}(\varepsilon)=\{R_{\theta}(\varepsilon):\theta\in K\}$. Following Haussler and Welzl (1986), the VC dimension of the range space $(K,\mathcal{R}(\varepsilon))$ is the largest integer $v$ such that $\mathcal{R}(\varepsilon)$ shatters some set $X\subset K$ of size $v$, where shattering means that for every subset $Y\subseteq X$ there exists a range $R\in\mathcal{R}(\varepsilon)$ such that $R\cap X=Y$. The following result is proved in Diakonikolas and Yannakakis (2010):
Lemma 4 (Lemma 4.5 in Diakonikolas and Yannakakis (2010)).
For any $\varepsilon>0$, the VC dimension of the range space $(K,\mathcal{R}(\varepsilon))$ is at most $d$.
Note that Lemma˜4 does not require $K$ to be convex. We now depart from the reduction in Diakonikolas and Yannakakis (2010) and exploit convexity. For each $\theta\in K$ set
$Q_{\theta}(\varepsilon)=\frac{1}{1+\varepsilon}\theta+\frac{\varepsilon}{1+\varepsilon}K.$
By convexity of $K$, we have $Q_{\theta}(\varepsilon)\subset K$ for every $\theta\in K$. Moreover, for every $\theta\in K$ it holds that
$Q_{\theta}(\varepsilon)\subset R_{\theta}(\varepsilon).$
Indeed, any $\phi\in Q_{\theta}(\varepsilon)$ can be written as $\phi=\frac{1}{1+\varepsilon}\theta+\frac{\varepsilon}{1+\varepsilon}\nu$ with $\nu\in K$, and then for all $j\in[d]$,
$\theta_{j}\leqslant(1+\varepsilon)\left(\frac{1}{1+\varepsilon}\theta_{j}+\frac{\varepsilon}{1+\varepsilon}\nu_{j}\right)=(1+\varepsilon)\phi_{j},$
so $\phi\in R_{\theta}(\varepsilon)$. Let $\mu$ be the uniform probability measure on $K$, which is well-defined since $\mathrm{Vol}(K)>0$. Then for any $\theta\in K$, a volumetric argument gives
$\mu(R_{\theta}(\varepsilon))\geqslant\mu(Q_{\theta}(\varepsilon))=\left(\frac{\varepsilon}{1+\varepsilon}\right)^{d}.$
Thus, it suffices to construct a set $A$ such that for all $\theta\in K$ one has $A\cap R_{\theta}(\varepsilon)\neq\emptyset$. Now the class $\mathcal{R}(\varepsilon)$ has VC dimension at most $d$, and every range satisfies the uniform lower bound $\mu(R_{\theta}(\varepsilon))\geqslant\eta$, where $\eta=\left(\frac{\varepsilon}{1+\varepsilon}\right)^{d}.$ Thus by a standard VC lower bound (see Haussler and Welzl (1986) and (Blumer et al., 1989, Theorem 2.1)), if we sample
$\left\lceil\frac{8d}{\eta}\log\left(\frac{13}{\eta}\right)\right\rceil$
points independently according to $\mu$ and let $A$ be the resulting set, then with probability at least $1/2$ we have $A\cap R_{\theta}(\varepsilon)\neq\emptyset$ for all $\theta\in K$. Plugging in the value of $\eta$ yields the stated bound. Existence follows since this event has positive probability. ∎
5 Mixture models (convex aggregation)
We are ready to state our second main statistical result.
Theorem 6.
Let $X_{1},\ldots,X_{n}$ with $n\geqslant 2$ be an i.i.d. sample from some unknown density $p^{\star}$ that belongs to the convex hull of a known class of arbitrary densities $\{p_{1},\ldots,p_{M}\}$, namely $\mathcal{P}=\operatorname{conv}\{p_{1},\ldots,p_{M}\}$ defined on the space $\mathcal{X}$ with respect to the measure $\mu$ and $p^{\star}\in\mathcal{P}$. There is a density estimator $\widehat{p}$ such that, with probability at least $1-\delta$,
$\operatorname{KL}\left(p^{\star},\widehat{p}\right)\leqslant\frac{870030\,M+45000\log(M)\log\left(4/\delta\right)}{n}.$
In the regime $n\geqslant\log(1/\delta)$, provided that $n\geqslant M\geqslant 5000$, this bound is optimal in the sense that there are $M$ distributions such that no estimator can perform better up to absolute constant factors, as shown in (Mourtada, 2025, Theorem 4).
As above, for the sake of presentation we assume that we are given a sample of size $2n$ (we translate back to the original sample size at the end). In general, provided that the sample size is at least $2$ we can split the sample into two equally sized independent parts, while dropping at most one element of the original sample. This will only lead to a larger absolute constant in the final statement. We describe the estimator $\widehat{p}$ of Theorem 6 as follows:
  • Use the first $n$ observations to run a Birgé-Le Cam tournament (see Lemma˜9) over $\mathcal{P}$ to find a convex closed set $\widehat{\mathcal{Q}}_{1}\subset\mathcal{P}$ such that, with probability $1-\delta/2$, it satisfies
    $\sup\limits_{p,q\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\widehat{\mathcal{Q}}_{1}.$ (13)
  • Let $\Delta$ be the simplex in $\mathbb{R}^{M}$. Consider the closed convex set $\widehat{\Delta}\subseteq\Delta$ defined as follows:
    $\widehat{\Delta}=\left\{\theta\in\Delta:\sum_{j=1}^{M}\theta_{j}p_{j}\in\widehat{\mathcal{Q}}_{1}\right\},$
    and construct a $32$-ratio cover of $\widehat{\Delta}$ with cardinality at most $2^{8M}$, denoted as $\widehat{\Delta}_{c}$.
  • Choose any $\beta^{(1)},\ldots,\beta^{(M)}\in\widehat{\Delta}$ such that for all $j\in[M]$, $\beta^{(j)}_{j}=\sup\{\theta_{j}:\theta\in\widehat{\Delta}\}$, i.e., each vector $\beta^{(j)}$ maximizes the $j$-th coordinate in $\widehat{\Delta}$. Construct a (random) density class
    $\widehat{\mathcal{Q}}_{2}=\left\{\frac{1}{2}\sum\limits_{j=1}^{M}\theta_{j}p_{j}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}:\theta\in\widehat{\Delta}_{c}\right\}.$ (14)
  • Using the second half of the sample, run the density estimator of Lemma˜2 over $\widehat{\mathcal{Q}}_{2}$ to output the final density $\widehat{p}$.
Proof of Theorem 6.
Fix $\delta\in(0,1)$. Combining the Birgé-Le Cam tournament result (Lemma˜9) and the local Hellinger entropy bound of Theorem˜2, we construct the desired convex set $\widehat{\mathcal{Q}}_{1}\subset\mathcal{P}$ satisfying, with probability at least $1-\delta/2$,
$\sup\limits_{p,q\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\widehat{\mathcal{Q}}_{1}.$ (15)
From now on we work under this event.
Using the second half of the sample, we output the density $\widehat{p}$ obtained using the procedure of Lemma˜2 with respect to $\widehat{\mathcal{Q}}_{2}$ defined by (14). Let $\theta\in\widehat{\Delta}$ be such that $p^{\star}=\sum\limits_{j=1}^{M}\theta_{j}p_{j}$. By definition of $\widehat{\Delta}_{c}$, there exists a vector $\phi\in\widehat{\Delta}_{c}$ such that
$\theta_{j}\leqslant 32\,\phi_{j}\quad\text{for every }j=1,\dots,M.$ (16)
We then define
$p^{\star}_{c}=\sum_{j=1}^{M}\phi_{j}p_{j}\quad\text{and}\quad\widetilde{p}=\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}.$
[…]
$\sup_{p_{\alpha},p_{\alpha^{\prime}}\in\widehat{\mathcal{Q}}_{2}}\sup_{x\in\mathcal{X}}\log\left(\frac{p_{\alpha}(x)}{p_{\alpha^{\prime}}(x)}\right)\leqslant\log\left(1+\frac{\sum_{k}\alpha_{k}p_{k}(x)}{\sum_{k}\bar{\beta}_{k}p_{k}(x)}\right)\leqslant\log\left(1+M\right)\leqslant\log(2M).$
Therefore, by Lemma˜2, and since by Theorem 3 it holds that $|\widehat{\mathcal{Q}}_{2}|\leqslant 256^{M}$, we have, with probability at least $1-\delta/2$,
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\operatorname{KL}({p^{\star}},{\widetilde{p}})+\frac{2M\log(256)+\frac{25}{4}\left(e-2\right)\left(\log(2M)+1\right)\log\left(\frac{4}{\delta}\right)}{n}.$ (17)
Finally, by Lemma 1 and the convexity of the Hellinger distance squared, and observing that the ratio of $p^{\star}$ to $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}$ is bounded by $64$ by (16), and, in particular, $p^{\star}(x)\leqslant 32\,p^{\star}_{c}(x).$ Therefore,
[…]
$\leqslant\frac{2\log(64)\left(3528\log(64)\,M+512\log(4/\delta)\right)}{(\sqrt{e}-1)^{2}\,n},$
where in the last line we used that $p^{\star}\in\widehat{\mathcal{Q}}_{1}$ and $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}\in\widehat{\mathcal{Q}}_{1}$ combined with (15). Plugging this into (17) and using the union bound, we have with probability at least $1-\delta$,
[…]
$\qquad+\frac{2\log(64)\left(3528\log(64)\,M+512\log\left(4/\delta\right)\right)}{(\sqrt{e}-1)^{2}n}.$
Note that we used a sample of size $2n$, which leads to an increased absolute constant by at most a factor of $3$. The claim follows.∎
6 Applications of ratio covers
In this section, we will discuss two more applications of our technique for ratio covering of convex sets developed in Section 4. In Subsection 6.1, we provide a sharpening (in some cases) of the result of Yang-Barron for obtaining expected error guarantees from KL covers, which will also yield our desired expected error guarantee for mixture models. In Subsection 6.2, we partially resolve a problem of Tang Tang (2022) about KL covers for the simplex via Theorem 3.
6.1 Expected error guarantees: sharpening the Yang–Barron bound
The celebrated result of Yang and Barron Yang and Barron (1999) shows how to turn a finite KL cover into an expected KL risk bound via a progressive mixture construction. In this section, we explain how our ratio-cover approach yields a sharper guarantee when the underlying density class is convex and we may construct appropriate ratio covers. We use the following form of the Yang–Barron bound.
Theorem 7 (Yang–Barron; see the proof of Theorem 32.1 in Section 32.2.1 of Polyanskiy and Wu (2025)).
Let $X_{1},\ldots,X_{n}$ be i.i.d. from a density $p^{\star}$ on $(\mathcal{X},\mu)$, and let $q_{1},\dots,q_{N}$ be densities such that
$\min_{j\in[N]}\operatorname{KL}(p^{\star},q_{j})\leqslant\varepsilon.$
Define the progressive mixture estimator $\widehat{p}=\sum_{j=1}^{N}w_{j}q_{j}$, where
$w_{j}=\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{\prod_{t=1}^{i-1}q_{j}(X_{t})}{\sum_{k=1}^{N}\prod_{t=1}^{i-1}q_{k}(X_{t})}.$
Then
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant\varepsilon+\frac{\log N}{n+1}.$
Two issues prevent a direct use of Theorem˜7 in our setting. First, constructing KL covers is substantially more delicate than constructing Hellinger covers. Second, even when one has sharp KL covering numbers (e.g., for the simplex), applying Yang–Barron with a global cover introduces unnecessary logarithmic factors, since it does not exploit localization.
[…]
Lemma 5.
Let $\mathcal{P}$ be a convex class of densities on $\mathcal{X}$ with respect to $\mu$, fix $p\in\mathcal{P}$, $\varepsilon>0$, and $\alpha\geqslant 2$. Set
$\mathcal{Q}=\mathcal{P}\cap B_{\operatorname{H}^{2}}(p,\varepsilon),\qquad B_{\operatorname{H}^{2}}(p,\varepsilon)=\{q\ \text{is a density}:\ \operatorname{H}^{2}(p,q)\leqslant\varepsilon\}.$
Assume that $S\subset\mathcal{Q}$ is finite and has the following property: for every $q\in\mathcal{Q}$ there exists $\widetilde{q}\in S$ such that
$q(x)\leqslant\alpha\,\widetilde{q}(x)\qquad\text{for all }x\in\mathcal{X}.$
Then for every $q\in\mathcal{P}$,
$\min_{\widetilde{q}\in S}\operatorname{KL}(q,\widetilde{q})\leqslant 30\,\bigl(\varepsilon+\operatorname{H}^{2}(p,q)\bigr)\,\log\Bigl(\alpha\Bigl(1\vee\frac{\operatorname{H}^{2}(p,q)}{\varepsilon}\Bigr)\Bigr).$
The proof of this result is deferred to Appendix˜D. Combined with the tail bounds on the Hellinger estimator, we obtain an expected KL guarantee for convex classes admitting suitable ratio covers.
Proposition 1.
Assume that $n$ is even. Let $\mathcal{P}$ be a convex class of densities and assume that $X_{1},\ldots,X_{n}$ be i.i.d. from some $p^{\star}\in\mathcal{P}$. Let $\varepsilon_{n/2}>0$ satisfy
$\frac{n}{2}\varepsilon_{n/2}^{2}\geqslant\log N_{\mathsf{loc}}(\mathcal{P},\varepsilon_{n/2})\,\lor\,1.$
Assume that there exist $\alpha\geqslant 2$ and an integer $N_{S}\geqslant 1$ such that for every $p\in\mathcal{P}$ one can find a finite set $S(p)\subset\mathcal{P}\cap B_{\operatorname{H}^{2}}(p,\varepsilon_{n/2}^{2})$ which is an $\alpha$-ratio cover of $\mathcal{P}\cap B_{\operatorname{H}^{2}}(p,\varepsilon_{n/2}^{2})$ (in the sense of Lemma˜5) and satisfies $|S(p)|\leqslant N_{S}$. Then, there exists an estimator $\widehat{p}$ such that
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 366609\,\varepsilon_{n/2}^{2}\log(\alpha)+\frac{\log(N_{S})}{n/2+1}.$
Proof.
Split the sample into two independent halves. Let $\widehat{p}_{\operatorname{H}}$ be the Le Cam–Birgé estimator (Lemma˜9) built from the first $n/2$ observations, and let $\widehat{p}$ be the Yang–Barron mixture (Theorem 7) built from the remaining $n/2$ observations over the candidate set $S(\widehat{p}_{\operatorname{H}})$. Conditioning on the first half and applying Theorem 7 to the second half yields
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant{\mathbf{E}}\Big[\min_{q\in S(\widehat{p}_{\operatorname{H}})}\operatorname{KL}(p^{\star},q)\Big]+\frac{\log(N_{S})}{n/2+1}.$
By Lemma 5 with $p=\widehat{p}_{\operatorname{H}}$ and $\varepsilon=\varepsilon_{n/2}^{2}$,
$\min_{q\in S(\widehat{p}_{\operatorname{H}})}\operatorname{KL}(p^{\star},q)\leqslant 30\Bigl(\varepsilon_{n/2}^{2}+\operatorname{H}^{2}(\widehat{p}_{\operatorname{H}},p^{\star})\Bigr)\log\Bigl(\alpha\Bigl(1\vee\frac{\operatorname{H}^{2}(\widehat{p}_{\operatorname{H}},p^{\star})}{\varepsilon_{n/2}^{2}}\Bigr)\Bigr).$
For $x\in(0,1)$, Lemma˜9 applied with sample size $n/2$ gives, with probability at least 1x1-x,
$\operatorname{H}^{2}(\widehat{p}_{\operatorname{H}},p^{\star})\leqslant 882\,\varepsilon_{n/2}^{2}+\frac{256\log(2/x)}{n}.$
Therefore,
[…]
Since $(n/2)\varepsilon_{n/2}^{2}\geqslant 1$, we have $\varepsilon_{n/2}^{2}\geqslant 2/n$, and hence
$\frac{256\log(2/x)}{\varepsilon_{n/2}^{2}\,n}\leqslant 128\log(2/x),\qquad\frac{256\log(2/x)}{n}\leqslant 128\,\varepsilon_{n/2}^{2}\log(2/x).$
Thus, we have
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 30\,\varepsilon_{n/2}^{2}\int_{0}^{1}\Bigl(883+128\log(2/x)\Bigr)\,\log\left(\alpha\Bigl(882+128\log(2/x)\Bigr)\right)\,dx+\frac{\log(N_{S})}{n/2+1}.$
Evaluating the integral yields
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 33000\,\varepsilon_{n/2}^{2}\log(\alpha)+231240\,\varepsilon_{n/2}^{2}+\frac{\log(N_{S})}{n/2+1}.$
Since $\alpha\geqslant 2$, we have $\log(\alpha)\geqslant\log 2$, so $231240\,\varepsilon_{n/2}^{2}\leqslant\frac{231240}{\log 2}\,\varepsilon_{n/2}^{2}\log(\alpha)$. This proves the claim.∎
When combined with the covering estimates of Theorem˜2 and Theorem˜3, this establishes the optimal in-expectation rate for mixture density estimation:
Corollary 3.
In the setup of Theorem˜6 let $n$ be even. There exists a density estimator $\widehat{p}$ such that
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 11000000\,\frac{M}{n}.$
Proof.
Apply Proposition 1 with $\mathcal{P}=\mathop{\mathrm{conv}}\{p_{1},\ldots,p_{M}\}$ and $\alpha=32$. By Theorem 2, $N_{\mathsf{loc}}(\mathcal{P},\varepsilon)\leqslant 64^{M-1}$ for all $\varepsilon>0$, so we may take
$\varepsilon_{n/2}^{2}=\frac{2M\log(64)}{n},$
which satisfies $(n/2)\varepsilon_{n/2}^{2}\geqslant\log N_{\mathsf{loc}}(\mathcal{P},\varepsilon_{n/2})$. Fix any $p\in\mathcal{P}$ and write $p=p_{v}$ for some $v\in\Delta_{M-1}$, where $p_{v}=\sum_{j=1}^{M}v_{j}p_{j}$. The set of mixture weights
$K=\Bigl\{u\in\Delta_{M-1}:\ p_{u}\in\mathcal{P}\cap B_{\operatorname{H}^{2}}(p,\varepsilon_{n/2}^{2})\Bigr\}$
is convex and compact. By Theorem 3, it admits a $32$-ratio cover of size at most $2^{8M}$; mapping these weights back to densities yields a $32$-ratio cover $S(p)$ of $\mathcal{P}\cap B_{\operatorname{H}^{2}}(p,\varepsilon_{n/2}^{2})$ with the same size. Hence Proposition 1 applies with $N_{S}=2^{8M}$. We obtain
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 366609\,\varepsilon_{n/2}^{2}\log(32)+\frac{8M\log 2}{n/2+1}.$
Using $1/(n/2+1)\leqslant 2/n$ and $\varepsilon_{n/2}^{2}=2M\log(64)/n$, the right-hand side is at most
$\frac{366609\cdot 2\log(64)\cdot\log(32)}{n}\,M+\frac{16\log 2}{n}\,M\leqslant\frac{10568316\,M}{n},$
which completes the proof.∎
6.2 KL covering numbers
Constructing covers in Kullback–Leibler divergence is substantially less developed than for metrics such as Hellinger distance. Tang’s PhD thesis Tang (2022) highlights the KL covering number of the simplex as a natural benchmark problem. This benchmark is directly tied to discrete distribution estimation: in this section, $\Delta_{d-1}$ is the convex hull of the $d$ Dirac masses on $[d]$, equivalently the class of mixtures of a finite dictionary of $d$ point masses.
We note that the results of this section are mainly illustrative. We focus on the simplex to mirror the setting studied in Tang (2022), but global KL covering numbers for the simplex are too crude for our statistical applications, where we crucially exploit localization. More broadly, our ratio-cover approach applies to convex hulls of arbitrary collections of densities, and our focus is always on local covering numbers.
[…]
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)=\min\Bigl\{N:\ \exists\,q^{(1)},\dots,q^{(N)}\in\Delta_{d-1}\ \text{such that}\ \sup_{p\in\Delta_{d-1}}\ \min_{k\in[N]}\ \operatorname{KL}({p},{q^{(k)}})\leqslant\varepsilon\Bigr\}.$
Tang’s KL covers for the simplex Tang (2022).
Tang gives a general lower bound (via a reduction to total variation covering numbers), which yields
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\geqslant\left(\frac{1}{8\varepsilon}\right)^{\frac{d-1}{2}}.$ (18)
On the other hand, for any $p\in\Delta_{d-1}$, the uniform distribution $u=(1/d,\dots,1/d)$ satisfies $\operatorname{KL}({p},{u})\leqslant\log d$, so $N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)=1$ whenever $\varepsilon\geqslant\log d$. For $0<\varepsilon<\log d$, Tang’s Theorem 2 gives the upper bound
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant\left(C\frac{\log d}{\varepsilon}\right)^{\frac{d-1}{2}}$
for a universal constant $C>0$, using a Hellinger cover together with a uniform lower bound on coordinates to control likelihood ratios. In the small-$\varepsilon$ regime $0<\varepsilon<\frac{1}{4(d+1)^{2}}$, Tang’s Theorem 3 gives the sharper bound
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant\left(\frac{C}{\varepsilon}\right)^{\frac{d-1}{2}}\log\frac{1}{\varepsilon},$
via a probabilistic construction, namely the proof uses a $\chi^{2}$-type argument based on a modification of Jeffreys prior. These results motivate the following question: for a fixed accuracy level (say $0<\varepsilon\leqslant 1$), can one match the lower bound (18) up to universal constants, i.e., obtain an upper bound of the form
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant\left(\frac{C}{\varepsilon}\right)^{\frac{d-1}{2}}\qquad(0<\varepsilon\leqslant 1)$
for some universal constant $C>0$? We answer this question in the affirmative using ratio covers (see Proposition˜2 below). We note that Tang also analyzes additional regimes where $\varepsilon$ varies with $d$, including subexponential and polynomial behaviors; for an overview we reproduce their phase diagram in Table 1.
Simplex KL Covering Number Bounds of Tang (2022)
Name Regime Lower Bound Upper Bound
Hellinger method $0<\varepsilon<\log d$ $\left(\dfrac{c}{\varepsilon}\right)^{\frac{d-1}{2}}$ $\left(\dfrac{C\log d}{\varepsilon}\right)^{\frac{d-1}{2}}$
$\chi^{2}$ method $0<\varepsilon<\dfrac{1}{4(d+1)^{2}}$ $\left(\dfrac{c}{\varepsilon}\right)^{\frac{d-1}{2}}$ $\log\left(\dfrac{1}{\varepsilon}\right)\left(\dfrac{C}{\varepsilon}\right)^{\frac{d-1}{2}}$
Subexponential $\varepsilon=\frac{1}{r}\log d+\log(r+1),\;1\leqslant r\leqslant e\log d$ $\exp\left(\frac{1}{e^{2}}d^{\,1-1/r}\right)$ $\exp({d^{\,1-1/r}\log d})$
Polynomial $\varepsilon=\log\frac{d}{\gamma},\qquad\gamma\leqslant\sqrt{\frac{d}{2}}$ $\exp({\frac{\gamma}{e}+\log\sqrt{\frac{2\pi\gamma}{e}}})$ $\exp(2\gamma\log d)$
Table 1: Covering number bounds for the simplex in Tang (2022) for different regimes of $\varepsilon$. Constants $c$ and $C$ may vary for each bound. This table corresponds to their Figure 3.2. Our ratio cover results will improve the regimes affected by the Hellinger method and the $\chi^{2}$ method.
Our next result improves the KL covering number upper bound in the regime where $\varepsilon$ is at most an absolute constant. We also state a general reduction from KL covers of the simplex to local Hellinger covers, combined with ratio covers.
Proposition 2.
Let $N_{\operatorname{H}^{2}}(\Delta_{d-1},r)$ denote the covering number of the simplex $\Delta_{d-1}$ in squared Hellinger distance at scale $r$. Then, for every $\varepsilon>0$,
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant N_{\operatorname{H}^{2}}\left(\Delta_{d-1},\frac{\varepsilon}{70}\right)\cdot 2^{8d}.$
In particular, for every $0<\varepsilon\leqslant 1$,
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant\left(\frac{5040\cdot 2^{32}}{\varepsilon}\right)^{\frac{d-1}{2}}.$
Proof.
If $d=1$, then the set is the singleton and the proof is immediate. Thus, we assume $d\geqslant 2$. Fix $\varepsilon>0$ and set $R=\varepsilon/70$. Let $\{q^{(1)},\dots,q^{(N)}\}\subset\Delta_{d-1}$ be an $R$-cover of $\Delta_{d-1}$ in squared Hellinger distance, so that for every $p\in\Delta_{d-1}$ there exists $k\in[N]$ with $\operatorname{H}^{2}(p,q^{(k)})\leqslant R$, and $N=N_{\operatorname{H}^{2}}(\Delta_{d-1},R)$. For each $k\in[N]$, define the (squared-Hellinger) ball intersection
$K_{k}=\Delta_{d-1}\cap B_{\operatorname{H}^{2}}(q^{(k)},R)\qquad\text{where}\qquad B_{\operatorname{H}^{2}}(q,R)=\{p\in\Delta_{d-1}:\ \operatorname{H}^{2}(p,q)\leqslant R\}.$
Applying Theorem˜3 to a convex and compact set $K_{k}$, we obtain a $32$-ratio cover $A_{k}\subset K_{k}$ with $|A_{k}|\leqslant 2^{8d}$. Define $\mathcal{A}=\bigcup_{k=1}^{N}A_{k}$, so
$|\mathcal{A}|\leqslant N_{\operatorname{H}^{2}}(\Delta_{d-1},R)\cdot 2^{8d}.$
Now fix $p\in\Delta_{d-1}$ and pick $k\in[N]$ such that $\operatorname{H}^{2}(p,q^{(k)})\leqslant R$. Since $p\in K_{k}$ and $A_{k}$ is a $32$-ratio cover of $K_{k}$, there exists $q^{\star}\in A_{k}$ such that $p_{i}\leqslant 32\,q^{\star}_{i}$ for all $i\in[d]$. In particular, whenever $p_{i}>0$ we have $q^{\star}_{i}>0$ and $\log(p_{i}/q^{\star}_{i})\leqslant\log 32$, hence $\max_{i\in[d]}\log\frac{p_{i}}{q^{\star}_{i}}\leqslant\log 32.$ Moreover, since $q^{\star}\in A_{k}\subset K_{k}\subset B_{\operatorname{H}^{2}}(q^{(k)},R)$, we also have $\operatorname{H}^{2}(q^{(k)},q^{\star})\leqslant R$. Using the triangle inequality for $\operatorname{H}$,
$\operatorname{H}(p,q^{\star})\leqslant\operatorname{H}\bigl(p,q^{(k)}\bigr)+\operatorname{H}\bigl(q^{(k)},q^{\star}\bigr)\leqslant\sqrt{R}+\sqrt{R}=2\sqrt{R},$
and thus $\operatorname{H}^{2}(p,q^{\star})=\operatorname{H}(p,q^{\star})^{2}\leqslant 4R$. By Lemma˜1,
$\operatorname{KL}({p},{q^{\star}})\leqslant\frac{2}{(\sqrt{e}-1)^{2}}\operatorname{H}^{2}(p,q^{\star})\max\left\{1,\max_{i\in[d]}\log\frac{p_{i}}{q^{\star}_{i}}\right\}\leqslant\frac{8}{(\sqrt{e}-1)^{2}}\max\{1,\log 32\}\,R.$
The numerical constant satisfies $\frac{8}{(\sqrt{e}-1)^{2}}\max\{1,\log 32\}<70$, so with $R=\varepsilon/70$ we have $\operatorname{KL}({p},{q^{\star}})\leqslant\varepsilon$. Since $p\in\Delta_{d-1}$ was arbitrary, $\mathcal{A}$ is a KL $\varepsilon$-cover of $\Delta_{d-1}$, and therefore
$N_{\operatorname{KL}}(\Delta_{d-1},\varepsilon)\leqslant N_{\operatorname{H}^{2}}\left(\Delta_{d-1},\frac{\varepsilon}{70}\right)\cdot 2^{8d}.$
It remains to derive the stated $(C/\varepsilon)^{(d-1)/2}$ bound when $0<\varepsilon\leqslant 1$. Consider the bijection $p\mapsto\sqrt{p}=(\sqrt{p_{1}},\dots,\sqrt{p_{d}})$ between $\Delta_{d-1}$ and the positive orthant of the unit sphere ${S}^{d-1}_{+}=\{x\in\mathbb{R}_{+}^{d}:\ \|x\|_{2}=1\}.$ For $p,q\in\Delta_{d-1}$, we have $\operatorname{H}^{2}(p,q)=\frac{1}{2}\bigl\|\sqrt{p}-\sqrt{q}\bigr\|_{2}^{2},$ which implies
$N_{\operatorname{H}^{2}}(\Delta_{d-1},r)=N_{\ell_{2}}\left({S}^{d-1}_{+},\sqrt{2r}\right)\leqslant N_{\ell_{2}}\left({S}^{d-1},\sqrt{2r}\right),$
where $N_{\ell_{2}}({S}^{d-1},t)$ denotes the $\ell_{2}$ covering number of the unit sphere at radius $t$. For $0<t\leqslant 1$, a standard shell-volumetric argument yields $N_{\ell_{2}}({S}^{d-1},t)\leqslant\left(\frac{12}{t}\right)^{d-1}.$ Fix $r=\varepsilon/70$. For $0<\varepsilon\leqslant 1$, we have $\sqrt{2r}=\sqrt{2\varepsilon/70}\leqslant 1$, hence
$N_{\operatorname{H}^{2}}\left(\Delta_{d-1},\frac{\varepsilon}{70}\right)\leqslant\left(\frac{5040}{\varepsilon}\right)^{\frac{d-1}{2}}.$
Combining this with the first part of the proof, we prove the claim. ∎
7 Aggregation in the misspecified case
In this section we discuss aggregation under misspecification, where we do not assume that $p^{\star}\in\mathcal{P}$. This problem is also of interest in the aggregation literature. The results in this section have the following features:
  • we do not assume that $p^{\star}\in\mathcal{P}$;
  • the algorithms are simpler, and their analysis relies on less sophisticated arguments;
  • however, their statistical guarantees feature additional polylogarithmic factors in $n$ and $M$.
In particular, obtaining optimal rates up to universal constant factors when $p^{\star}\notin\mathcal{P}$ remains an open question.
Proposition 3.
Let $X_{1},\ldots,X_{n}$ be an i.i.d. sample from some unknown density $p^{\star}$ defined on the space $\mathcal{X}$ with respect to the measure $\mu$. Let $\mathcal{P}=\{p_{1},\ldots,p_{M}\}$ be a set of densities with respect to $\mu$ defined on $\mathcal{X}$. There is a density estimator $\widehat{p}$ such that for any $\delta\in(0,1/2]$, with probability at least $1-\delta$,
$\operatorname{KL}\left(p^{\star},\widehat{p}\right)\leqslant\min_{j\in[M]}\operatorname{KL}\left(p^{\star},p_{j}\right)+\frac{19\log(2nM)\log(1/\delta)}{n}.$
Compared to Theorem˜1 the bound contains an additional dependence on $n$ under the logarithm. The estimator and the analysis are presented in Appendix˜D. The idea behind the estimator is to run the predictor of Lemma˜2 on a deterministically smoothed class $\mathcal{Q}$ obtained from the given set $\mathcal{P}$.
Our next result is an analog of Theorem˜6 in the misspecified case.
Proposition 4.
Let $X_{1},\ldots,X_{n}$ with $n\geqslant 2$ be an i.i.d. sample from some unknown density $p^{\star}$. Let $\mathcal{P}=\{p_{1},\ldots,p_{M}\}$ be a set of densities with respect to $\mu$ defined on $\mathcal{X}$. There is a density estimator $\widehat{p}$ such that, with probability at least $1-\delta$,
$\operatorname{KL}\left(p^{\star},\widehat{p}\right)\leqslant\inf\limits_{p\in\operatorname{conv}(\mathcal{P})}\operatorname{KL}\left(p^{\star},p\right)+\frac{43M\log^{2}(3nM)+22\log(3nM)\log(1/\delta)}{n}.$
The main difference with Theorem˜6 consists in additional squared logarithmic terms in the risk bound. The estimator in Proposition˜4 corresponds to the MLE over a smoothed density class:
  • Construct a class of densities
    $\mathcal{Q}=\bigg\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\operatorname{conv}(\mathcal{P})\bigg\}.$
  • Let $\widehat{p}$ be the MLE in $\mathcal{Q}$. That is,
    $\widehat{p}=\arg\min_{p\in\mathcal{Q}}\widehat{L}(p).$
The proof of Proposition˜4 follows relatively standard lines using techniques from empirical process theory and we defer it to Appendix˜D.
8 Acknowledgments
S.C. is supported by the NDSEG Fellowship Program, Tselil Schramm’s NSF CAREER Grant no. 2143246, and Gregory Valiant’s Simons Foundation Investigator Award and NSF award AF-2341890. J.M. is supported by a grant of the French National Research Agency (ANR), “Investissements d’Avenir” (LabEx Ecodec/ANR-11-LABX-0047).
Appendix A High-probability online-to-batch conversion
We present a self-contained version of the argument of Theorem 1 in [56], adapted to the logarithmic loss and a finite class of densities.
Consider $M$ distributions $p_{1},\ldots,p_{M}$ with corresponding densities $p_{1},\ldots,p_{M}$. Given the sample $X_{1},\ldots,X_{n}$ define recursively the following sequence of densities:
$\widehat{p}_{1}=\frac{1}{M}\sum\limits_{i=1}^{M}p_{i},$
and for $j\in\{1,2,\ldots,n-1\}$ set
$\widehat{p}_{j+1}=\sum\limits_{i=1}^{M}p_{i}\frac{\prod_{k=1}^{j}\left(\frac{1}{2}p_{i}(X_{k})+\frac{1}{2}\widehat{p}_{k}(X_{k})\right)}{\sum\nolimits_{h=1}^{M}\prod_{k=1}^{j}\left(\frac{1}{2}p_{h}(X_{k})+\frac{1}{2}\widehat{p}_{k}(X_{k})\right)},$ (19)
The final predictor will be given by
$\widehat{p}=\frac{1}{n}\sum\limits_{i=1}^{n}\widehat{p}_{i},$ (20)
which is a version of the classical progressive Bayesian mixture estimator [7, 63] with the key difference that in (20) for each individual weight in (19) we consider the product of shifted elements $\left(\frac{1}{2}p_{i}(X_{k})+\frac{1}{2}\widehat{p}_{k}(X_{k})\right)$ instead of the canonical product of $p_{i}(X_{k})$ in the original formulation.
We begin by providing the following lemma which provides the regret bound for the sequence $\widehat{p}_{1},\ldots,\widehat{p}_{n}$.
Lemma 6.
Let $p_{1},\ldots,p_{M}$ be densities and let $X_{1},\ldots,X_{n}$ be any sequence of points. For the sequence $(\widehat{p}_{t})_{t=1}^{n}$ defined by (19) and (20), it holds that
$\sum_{t=1}^{n}-\log\big(\widehat{p}_{t}(X_{t})\big)\leqslant\min_{j\in[M]}\sum_{t=1}^{n}-\log\left(\tfrac{1}{2}\widehat{p}_{t}(X_{t})+\tfrac{1}{2}p_{j}(X_{t})\right)+\log M.$ (21)
Proof.
The proof repeats the standard lines for the logarithmic loss [59]. See the proof of the more general Proposition 1 in [56]. ∎
We also need the following elementary result.
Lemma 7.
Let $m>0$ and $x,y\in\mathbb{R}^{+}$ such that $\left|\log\left(x/y\right)\right|\leqslant m$. Then,
$\log\left(\frac{x+y}{2}\right)\leqslant\frac{-\log x-\log y}{2}-\frac{\left(\log\left(x/y\right)\right)^{2}}{8(m+1)}.$ (22)
Proof.
Let $h(w)=-\log w$. Then $h^{\prime}(w)=-w^{-1}$ and $h^{\prime\prime}(w)=w^{-2}=(h^{\prime}(w))^{2}$. Fix $(x,y)\in\mathbb{R}^{2}$ and set $s=\frac{x+y}{2}$. Define
$g(w)=h(w)-h(s)-\frac{\left(h(w)-h(s)\right)^{2}}{\gamma}.$
A direct computation gives
$g^{\prime\prime}(w)=h^{\prime\prime}(w)\left(1-\frac{2}{\gamma}\left(h(w)-h(s)+1\right)\right).$
Since $\left|\log\left(x/y\right)\right|\leqslant m$, we have $\left|h(x)-h(y)\right|\leqslant m$. For any $w$ between $x$ and $y$, the monotonicity of $h$ implies $h(w)$ and $h(s)$ lie between $h(x)$ and $h(y)$, hence $h(w)-h(s)\leqslant\left|h(x)-h(y)\right|\leqslant m$. With $\gamma=2(m+1)$ we obtain $g^{\prime\prime}(w)\geqslant 0$ on the segment between $x$ and $y$. By convexity, $g(s)\leqslant\frac{1}{2}g(x)+\frac{1}{2}g(y)$. Since $g(s)=0$, this yields
$h(s)\leqslant\frac{h(x)+h(y)}{2}-\frac{\left(h(x)-h(s)\right)^{2}+\left(h(y)-h(s)\right)^{2}}{2\gamma}.$
Using $\left(a-c\right)^{2}+\left(b-c\right)^{2}\geqslant\frac{1}{2}\left(a-b\right)^{2}$ with $a=h(x)$, $b=h(y)$, $c=h(s)$, we get
$h\left(\frac{x+y}{2}\right)\leqslant\frac{h(x)+h(y)}{2}-\frac{\left(h(x)-h(y)\right)^{2}}{4\gamma}.$
Substituting $h(w)=-\log w$ and $\gamma=2(m+1)$ proves the claim. ∎
Another useful lemma is a simple version of Freedman’s inequality. For the version with these explicit constants we refer to [10, Theorem 1]. Throughout this section we denote ${\mathbf{E}}_{i-1}[\cdot]={\mathbf{E}}[\cdot|X_{1},\ldots,X_{i-1}]$.
Lemma 8.
Let $X_{1},\ldots,X_{n}$ be a martingale difference sequence adapted to a filtration $(\mathcal{F}_{i})_{i\leqslant n}$. Suppose that $|X_{i}|\leqslant m$ almost surely. Then for any $\delta\in(0,1)$ and $\lambda\in[0,1/m]$, with probability at least $1-\delta$,
$\sum\limits_{i=1}^{n}X_{i}\leqslant\lambda(e-2)\sum\limits_{i=1}^{n}{\mathbf{E}}_{i-1}[X_{i}^{2}]+\frac{\log(1/\delta)}{\lambda}.$
Finally, we restate Lemma˜2 and specify the predictor $\widehat{p}$.
Theorem 8.
Assume $X_{1},\ldots,X_{n}$ are i.i.d. random variables with arbitrary and unknown density $p^{\star}$ on $(\mathcal{X},\mu)$. Let $p_{1},\ldots,p_{M}$ be densities such that for some $m>0$,
$\sup_{x\in\mathcal{X}}\max_{i,j\in[M]}\left|\log\left(\frac{p_{i}(x)}{p_{j}(x)}\right)\right|\leqslant m.$
Let $\left(\widehat{p}_{i}\right)_{i=1}^{n}$ and $\widehat{p}=\frac{1}{n}\sum_{i=1}^{n}\widehat{p}_{i}$ be defined by (19) and (20). Then for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\min_{j\in[M]}\operatorname{KL}({p^{\star}},{p_{j}})+\frac{2\log\left(M\right)+\frac{25}{4}\left(e-2\right)\left(m+1\right)\log\left(\frac{1}{\delta}\right)}{n}.$
Proof.
Fix $j\in[M]$. Write $\ell_{i}(q)=-\log\left(q\left(X_{i}\right)\right)$ and $r_{i}=\ell_{i}\left(\widehat{p}_{i}\right)-\ell_{i}\left(p_{j}\right)=\log\left(\frac{p_{j}\left(X_{i}\right)}{\widehat{p}_{i}\left(X_{i}\right)}\right)$. Since $\widehat{p}_{i}\left(x\right)$ is a convex combination of $\{p_{k}\left(x\right)\}_{k=1}^{M}$ and $\left|\log\left(p_{a}\left(x\right)/p_{b}\left(x\right)\right)\right|\leqslant m$ for all $a,b\in[M]$, one has $\exp(-m)\leqslant\widehat{p}_{i}\left(x\right)/p_{j}\left(x\right)\leqslant\exp(m)$ and hence $\left|r_{i}\right|\leqslant m$ almost surely. Set
$v_{i}=\frac{r_{i}^{2}}{4\left(m+1\right)},\quad\textrm{and}\quad Z_{i}=r_{i}+v_{i}.$
By construction, $\widehat{p}_{i}$ is $\mathcal{F}_{i-1}$-measurable and $X_{i}\stackrel{{\scriptstyle d}}{{=}}X$ is independent of $\mathcal{F}_{i-1}$. By Jensen’s inequality,
${\mathbf{E}}\log\left(\widehat{p}(X)\right)+{\mathbf{E}}\log\left(p_{j}(X)\right)\leqslant\frac{1}{n}\sum_{i=1}^{n}{\mathbf{E}}_{i-1}\left[r_{i}\right].$ (23)
Apply Lemma 7 with $x=\widehat{p}_{i}\left(X_{i}\right)$ and $y=p_{j}\left(X_{i}\right)$ we get
$\log\left(\frac{\widehat{p}_{i}\left(X_{i}\right)+p_{j}\left(X_{i}\right)}{2}\right)\leqslant\frac{\ell_{i}\left(\widehat{p}_{i}\right)+\ell_{i}\left(p_{j}\right)}{2}-\frac{r_{i}^{2}}{8\left(m+1\right)}.$
Equivalently,
$r_{i}\leqslant 2\left(\widetilde{\ell}_{i}\left(\widehat{p}_{i}\right)-\widetilde{\ell}_{i}\left(p_{j}\right)\right)-\frac{r_{i}^{2}}{4\left(m+1\right)}=2\left(\widetilde{\ell}_{i}\left(\widehat{p}_{i}\right)-\widetilde{\ell}_{i}\left(p_{j}\right)\right)-v_{i},$ (24)
where $\widetilde{\ell}_{i}\left(q\right)=-\log\left(\frac{\widehat{p}_{i}\left(X_{i}\right)+q\left(X_{i}\right)}{2}\right)$. Summing (24) over $i$ and using Lemma 6 (which, for each $j$, gives $\sum_{i=1}^{n}\left(\widetilde{\ell}_{i}\left(\widehat{p}_{i}\right)-\widetilde{\ell}_{i}\left(p_{j}\right)\right)\leqslant\log\left(M\right)$) yields
$\sum_{i=1}^{n}r_{i}\leqslant 2\log\left(M\right)-\sum_{i=1}^{n}v_{i}.$ (25)
Consider the martingale differences $Y_{i}={\mathbf{E}}_{i-1}\left[Z_{i}\right]-Z_{i}$ with respect to $\left(\mathcal{F}_{i}\right)_{i=1}^{n}$. Since $\left|r_{i}\right|\leqslant m$, one has $\left|Z_{i}\right|\leqslant\left|r_{i}\right|+\frac{r_{i}^{2}}{4\left(m+1\right)}\leqslant m+\frac{m^{2}}{4\left(m+1\right)}$. Denote $\widetilde{m}=m+\frac{m^{2}}{4\left(m+1\right)}$. We have $\left|Y_{i}\right|\leqslant 2\widetilde{m}$ almost surely. Moreover, since $|r_{i}|\leqslant m$ we have $Z_{i}=r_{i}+\frac{1}{4(m+1)}r_{i}^{2}=r_{i}\Bigl(1+\frac{1}{4(m+1)}r_{i}\Bigr),$ which implies $Z_{i}^{2}=r_{i}^{2}\Bigl(1+\frac{1}{4(m+1)}r_{i}\Bigr)^{2}\leqslant r_{i}^{2}\Bigl(1+\frac{1}{4(m+1)}m\Bigr)^{2}=c_{m}^{2}r_{i}^{2},$ where $c_{m}=\frac{5m+4}{4(m+1)}$. Applying Lemma 8 with
$\lambda=\frac{1}{4\left(m+1\right)\left(e-2\right)c_{m}^{2}}\leqslant\frac{1}{2\widetilde{m}}$,
we obtain, with probability at least $1-\delta$,
$\sum_{i=1}^{n}\left({\mathbf{E}}_{i-1}\left[Z_{i}\right]-Z_{i}\right)\leqslant\lambda\left(e-2\right)\sum_{i=1}^{n}{\mathbf{E}}_{i-1}\left[Y_{i}^{2}\right]+\frac{\log\left(1/\delta\right)}{\lambda}\leqslant\sum_{i=1}^{n}{\mathbf{E}}_{i-1}\left[v_{i}\right]+\frac{\log\left(1/\delta\right)}{\lambda},$
where in the last inequality we used ${\mathbf{E}}_{i-1}[Y_{i}^{2}]\leqslant{\mathbf{E}}_{i-1}[Z_{i}^{2}]\leqslant 4(m+1)c_{m}^{2}{\mathbf{E}}_{i-1}[v_{i}]$. Expanding $Z_{i}=r_{i}+v_{i}$ and canceling $\sum_{i}{\mathbf{E}}_{i-1}\left[v_{i}\right]$ gives
$\sum_{i=1}^{n}{\mathbf{E}}_{i-1}[r_{i}]\leqslant\sum_{i=1}^{n}r_{i}+\sum_{i=1}^{n}v_{i}+\frac{\log\left(1/\delta\right)}{\lambda}.$ (26)
Combining (26) with (25) yields
$\sum_{i=1}^{n}{\mathbf{E}}_{i-1}[r_{i}]\leqslant 2\log\left(M\right)+\frac{\log\left(1/\delta\right)}{\lambda}.$
Inserting this into (23) and finally choosing $j\in[M]$ that minimizes $\left(-{\mathbf{E}}\log\left(p_{j}(X)\right)\right)$, we have
${\mathbf{E}}\log\left(\widehat{p}(X)\right)-\min_{j\in[M]}\left(-{\mathbf{E}}\log\left(p_{j}(X)\right)\right)\leqslant\frac{2\log\left(M\right)}{n}+\frac{\log\left(1/\delta\right)}{n\,\lambda}.$
Finally, $1/\lambda=4\left(m+1\right)\left(e-2\right)c_{m}^{2}=\frac{\left(e-2\right)\left(5m+4\right)^{2}}{4\left(m+1\right)}\leqslant\frac{25}{4}\left(e-2\right)\left(m+1\right)$, which gives (7). The claim follows. ∎
Appendix B Suboptimality of MLE and Bayesian model averaging
In this section, we show that simple baseline methods for model aggregation, such as MLE and Bayes model averaging, fail to achieve optimal guarantees in the distribution-free setting.
The following simple fact, mentioned in the introduction, shows that in the distribution-free setting, no selection procedure (such as MLE) achieves any nontrivial guarantee.
Fact 1.
For every $n\geqslant 1$, there exists a dictionary $\mathcal{P}=\{p_{1},p_{2}\}$ such that the following holds. For every estimator $\widehat{p}=\widehat{p}_{n}$ based on an i.i.d. sample $X_{1},\dots,X_{n}$ of size $n$ and taking values inside $\mathcal{P}$, there exists $p^{\star}\in\mathcal{P}$ such that if $X_{1},\dots,X_{n}\sim p^{\star}$, then $\mathbf{P}(\operatorname{KL}({p^{\star}},{\widehat{p}_{n}})=+\infty)\geqslant 0.99$.
Proof.
Let $\mathcal{X}=\{0,1,2\}$ and $\eta>0$ be such that $(1-\eta)^{n}=0.99$. Define the densities $p_{1},p_{2}$ on $\mathcal{X}$ (with respect to the counting measure) by $p_{1}(0)=1-\eta,p_{1}(1)=\eta,p_{1}(2)=0$ and $p_{2}(0)=1-\eta,p_{2}(1)=0,p_{2}(2)=\eta$. Let $E$ denote the event $\{X_{1}=\dots=X_{n}=0\}$ and $q$ the value taken by the estimator $\widehat{p}_{n}$ under $E$, such that (for both $p^{\star}=p_{1}$ and $p^{\star}=p_{2}$) $\mathbf{P}(\widehat{p}_{n}=q)\geqslant\mathbf{P}(E)=(1-\eta)^{n}=0.99$. Now if $q=p_{1}$ we take $p^{\star}=p_{2}$, and if $q=p_{2}$ we take $p^{\star}=p_{1}$; in both cases, $\operatorname{KL}({p^{\star}},{q})=+\infty$ and thus $\mathbf{P}(\operatorname{KL}({p^{\star}},{\widehat{p}_{n}})=+\infty)\geqslant 0.99$. ∎
Next, in the convex aggregation setting, Theorem 2 of [47] shows that any estimator that does not depend on $\delta$ must incur a suboptimal high-probability error bound. In particular, there exist instances (already with $M=2$) for which any expectation-optimal estimator, including the progressive mixture estimator of Yang and Barron (formally defined in Theorem˜7), incurs a lower bound of order
$\frac{\log(1/\delta)\log\log(1/\delta)}{n},$
which is worse than what is achieved by our $\delta$-dependent estimator in Theorem˜6.
In the remainder of this section, we show that an even stronger lower bound holds for Bayesian model averaging, which is arguably the most natural model aggregation procedure. For completeness, we also provide a matching upper bound for the case of two densities, showing that our lower bound is tight up to absolute constants. Our lower bound focuses on two Bernoulli distributions, as in the lower-bound construction of [47].
Our next result shows that Bayesian model averaging, while satisfying nontrivial risk bounds, fails to achieve the optimal dependence on $\delta$ even when $M=2$. In particular, they necessarily incur a $\log^{2}(1/\delta)$ factor instead of the optimal $\log(1/\delta)$ factor.
Proposition 5.
Let $p^{\star},q$ be two densities on $\mathcal{X}$ with respect to $\mu$, and let $X_{1},\ldots,X_{n}$ be i.i.d. from $p^{\star}$. Consider Bayesian model averaging with a uniform prior, which outputs
$\widehat{p}=(1-\widehat{\alpha})p^{\star}+\widehat{\alpha}\,q,\qquad\widehat{\alpha}=\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})+\prod_{i=1}^{n}q(X_{i})}.$
Then for any $\delta\in(0,1/2)$, with probability at least $1-\delta$,
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant\min\left\{\frac{10\log^{2}(2/\delta)}{n},5\log\left(\frac{2}{\delta}\right)\right\}.$
Moreover, taking $\mathcal{X}=\{0,1\}$, letting $q$ be the point mass at $1$, and letting $p^{\star}$ be the Bernoulli law with $(1-p^{\star}(0))^{n}=\delta$, we have, for any $\delta\in(0,e^{-2}]$ and with probability at least $\delta$,
$\operatorname{KL}(p^{\star},\widehat{p})\geqslant\min\left\{\frac{\log^{2}(1/\delta)}{4n},\,\frac{1}{4}\log\left(\frac{1}{\delta}\right)\right\}.$
Proof.
By Markov’s inequality and ${\mathbf{E}}\big[\prod_{i=1}^{n}q(X_{i})/p^{\star}(X_{i})\big]\leqslant 1$,
$\mathbf{P}\left(\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}\geqslant\frac{1}{\delta}\right)\leqslant\delta.$
Hence, on an event of probability at least $1-\delta$ we have $\frac{\prod q(X_{i})}{\prod p^{\star}(X_{i})}<1/\delta$, which implies
$1-\widehat{\alpha}=\frac{1}{1+\frac{\prod q(X_{i})}{\prod p^{\star}(X_{i})}}\geqslant\frac{\delta}{1+\delta}\geqslant\frac{\delta}{2}.$
Therefore, for all $x$,
$\widehat{p}(x)=(1-\widehat{\alpha})p^{\star}(x)+\widehat{\alpha}q(x)\geqslant(1-\widehat{\alpha})p^{\star}(x)\geqslant\frac{\delta}{2}p^{\star}(x),$
and so
$\sup_{x}\log\left(\frac{p^{\star}(x)}{\widehat{p}(x)}\right)\leqslant\log\left(\frac{2}{\delta}\right).$ (27)
By Lemma 1,
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\operatorname{H}^{2}(p^{\star},\widehat{p})\log\left(\frac{2}{\delta}\right).$ (28)
We consider the cases. First, if $2\log(2/\delta)/n>1$, then the minimum in the statement is achieved at the second term and since $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant 1$, we have
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\log\left(\frac{2}{\delta}\right),$
which proves the claim. Otherwise, we have $2\log(2/\delta)/n\leqslant 1$, and we again consider two cases. In the first case, assume $\operatorname{H}^{2}(p^{\star},q)\leqslant\frac{2\log(2/\delta)}{n}$. Using (27) and the convexity of the squared Hellinger distance in its second argument, $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant\operatorname{H}^{2}(p^{\star},q)$, which implies
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant\frac{10\log^{2}(2/\delta)}{n}.$
Otherwise, we have $\operatorname{H}^{2}(p^{\star},q)>\frac{2\log(2/\delta)}{n}$. In this case we want to upper bound $\widehat{\alpha}$. We have as in the proof of Lemma˜3 for any $t>0$,
$\mathbf{P}\left(\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}\geqslant t\right)\leqslant\frac{1}{\sqrt{t}}\bigl(1-\operatorname{H}^{2}(p^{\star},q)\bigr)^{n}\leqslant\frac{1}{\sqrt{t}}\exp\bigl(-n\operatorname{H}^{2}(p^{\star},q)\bigr).$
We choose $t=\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)}$. Note that $t<1<1/\delta$. By the above line we have
$\mathbf{P}\left(\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}\geqslant\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)}\right)\leqslant\sqrt{\frac{n\operatorname{H}^{2}(p^{\star},q)}{2\log(2/\delta)}}\exp(-n\operatorname{H}^{2}(p^{\star},q))\leqslant\exp(-2\log(2/\delta))\leqslant\delta,$
where we used that $\sqrt{x}\exp(-x)$ is decreasing for $x\geqslant 1/2$ and $n\operatorname{H}^{2}(p^{\star},q)>2\log(2/\delta)>1/2$. Therefore, on the complementary event, $\widehat{\alpha}\leqslant\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}<\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)},$ and (27) also holds since $t<1/\delta$. Hence, by (28) and the convexity of the squared Hellinger distance in its second argument, we have on this event
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\operatorname{H}^{2}(p^{\star},\widehat{p})\log\left(\frac{2}{\delta}\right)\leqslant 5\widehat{\alpha}\operatorname{H}^{2}(p^{\star},q)\log\left(\frac{2}{\delta}\right)\leqslant\frac{10\log^{2}(2/\delta)}{n}.$
The upper bound follows.
For the lower bound, let $\mathcal{X}=\{0,1\}$. Let $q$ be the point mass at $1$ (so $q(1)=1,q(0)=0$). Fix $p\in(0,1)$ such that $(1-p)^{n}=\delta$, and let $p^{\star}$ be the Bernoulli law with $p^{\star}(0)=p$ and $p^{\star}(1)=1-p$. On the event
$A=\{X_{1}=\cdots=X_{n}=1\},$
which satisfies $\mathbf{P}(A)=(1-p)^{n}=\delta$, we have
$\widehat{\alpha}=\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})+\prod_{i=1}^{n}q(X_{i})}=\frac{1}{1+\delta},\qquad\widehat{p}(0)=(1-\widehat{\alpha})p=\frac{p\delta}{1+\delta}.$
By the Bernoulli KL formula,
$\operatorname{KL}(p^{\star},\widehat{p})=(1-p)\log\left(\frac{(1+\delta)(1-p)}{1+\delta-p\delta}\right)+p\log\left(\frac{1+\delta}{\delta}\right).$
Since $1+\delta-p\delta\leqslant 1+\delta$, the ratio $\frac{(1+\delta)(1-p)}{1+\delta-p\delta}$ lies in [1p,1][1-p,1]. Hence
$(1-p)\log\left(\frac{(1+\delta)(1-p)}{1+\delta-p\delta}\right)\geqslant(1-p)\log(1-p)\geqslant-p,$
where the last inequality follows from $\log(1-x)\leqslant\frac{x}{1-x}$. Therefore, on $A$,
$\operatorname{KL}(p^{\star},\widehat{p})\geqslant p\log\left(\frac{1+\delta}{\delta}\right)-p\geqslant p\log\left(\frac{1}{\delta}\right)-p.$ (29)
We now show that $p\geqslant\frac{\log(1/\delta)}{n+\log(1/\delta)}$. Using $1-x\geqslant\exp\left(-\frac{x}{1-x}\right)$ for $x\in(0,1)$, we have
$\delta=(1-p)^{n}\geqslant\exp\left(-\frac{pn}{1-p}\right),$
which implies $p\geqslant\frac{\log(1/\delta)}{n+\log(1/\delta)}$. Thus, if $\log(1/\delta)\leqslant n$ then $p\geqslant\frac{1}{2}\frac{\log(1/\delta)}{n}$. Since $\delta\leqslant e^{-2}$ implies $\log(1/\delta)\geqslant 2$, from (29) we get
$\operatorname{KL}(p^{\star},\widehat{p})\geqslant p\bigl(\log(1/\delta)-1\bigr)\geqslant\frac{1}{2}\,p\log\left(\frac{1}{\delta}\right).$
Therefore, for $\delta\in(0,e^{-2}]$, it holds that
$\operatorname{KL}(p^{\star},\widehat{p})\ \geqslant\ \frac{1}{2}\,p\log\left(\frac{1}{\delta}\right)\ \geqslant\ \begin{cases}\displaystyle\frac{1}{4}\frac{\log^{2}(1/\delta)}{n},&\text{if }\ \log(1/\delta)\leqslant n,\\[5.16663pt] \displaystyle\frac{1}{4}\log\left(\frac{1}{\delta}\right),&\text{if }\ \log(1/\delta)\geqslant n,\end{cases}$
where, in the second case, we also used $p\geqslant\frac{\log(1/\delta)}{n+\log(1/\delta)}\geqslant\frac{1}{2}$. Since $\mathbf{P}(A)=\delta$, these lower bounds hold with probability at least $\delta$. This completes the proof. ∎
Appendix C Birgé–Le Cam tournament
In this section we reproduce the details of the Birgé–Le Cam tournament procedure. The derivations are standard and we mainly follow the exposition in [50, Section 32.2] with the minor difference that we need to provide sharper tail bounds for our applications. For earlier results establishing this procedure we refer to [38, 14, 15]. For a set of densities $\mathcal{P}$, recall the definition (10) of the local Hellinger entropy. Let $\varepsilon_{n}>0$ satisfy the fixed-point inequality
$n\varepsilon_{n}^{2}\geqslant\log N_{\mathsf{loc}}(\mathcal{P},\varepsilon_{n})\lor 1.$ (30)
Lemma 9 (Le Cam-Birgé tournament).
Assume we observe an i.i.d. sample $X_{1},\ldots,X_{n}$ distributed according to the density $p^{\star}$ with respect to the measure $\mu$. Let $\mathcal{P}$ be a known class of densities such that $p^{\star}\in\mathcal{P}$. There exists an estimator $\widehat{p}=\widehat{p}(X_{1},\ldots,X_{n})\in\mathcal{P}$ such that for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\operatorname{H}(p^{\star},\widehat{p})\leqslant 21\varepsilon_{n}+8\sqrt{\frac{\log(2/\delta)}{n}},$
where $\varepsilon_{n}$ is given by (30). Moreover,
${\mathbf{E}}\operatorname{H}(p^{\star},\widehat{p})\leqslant 21\varepsilon_{n}+\frac{11}{\sqrt{n}}.$
We now present a construction of the estimator as well as the construction of the convex set containing $p^{\star}$, which will be used in our analysis.
  1. 1.
    Take a maximal $\varepsilon$-packing $\{p_{1},\ldots,p_{N}\}\subset\mathcal{P}$ in Hellinger distance $\operatorname{H}$.
  2. 2.
    For each pair $i\neq j$ with $\operatorname{H}(p_{i},p_{j})\geqslant R$ where $R=4\varepsilon$, run the composite test $\psi_{ij}$ of Lemma 10 between the hypotheses $B_{\operatorname{H}}(p_{i},\varepsilon)$ and $B_{\operatorname{H}}(p_{j},\varepsilon)$ and set $\psi_{ji}=1-\psi_{ij}$.
  3. 3.
    Define tournament scores
    $T_{i}=\begin{cases}\max\limits_{j:\ \operatorname{H}(p_{i},p_{j})\geqslant R,\ \psi_{ij}=1}\ \operatorname{H}(p_{i},p_{j}),&\text{if such }j\text{ exists},\\ 0,&\text{otherwise.}\end{cases}$
  4. 4.
    Let $i^{\star}\in\arg\min_{i}T_{i}$ (ties arbitrary) and output $\widehat{p}=p_{i^{\star}}$. Namely, we choose the density for which no $R$-distant center “wins the match” defined by the composite test $\psi_{ij}$.
  5. 5.
    For any $\delta\in(0,1)$ define the set
    $\widehat{\mathcal{Q}}_{1}=\left\{q\in\mathcal{P}:\operatorname{H}(q,\widehat{p})\leqslant 21\varepsilon_{n}+8\sqrt{\frac{\log(2/\delta)}{n}}\right\}.$
[…]
We first recall the following result on hypotheses testing between Hellinger balls, due to Le Cam [38] and Birgé [14].
Lemma 10 (Pairwise composite testing; equation (32.33) of [50]).
Let $\varepsilon>0$ and $R=4\varepsilon$. For any $i\neq j$ with $\operatorname{H}(p_{i},p_{j})\geqslant R$, there exists a test $\psi_{ij}(X_{1},\ldots,X_{n})\in\{0,1\}$ such that the outcome 0 means $p\in B_{\operatorname{H}}(p_{i},\varepsilon)$, the outcome $1$ means $p\in B_{\operatorname{H}}(p_{j},\varepsilon)$, and $\psi_{ji}=1-\psi_{ij}$. Moreover,
$\sup_{p\in B_{\operatorname{H}}(p_{i},\varepsilon)}\mathbf{P}_{p}[\psi_{ij}(X_{1},\ldots,X_{n})=1]\leqslant\exp\Big(-\frac{n}{4}\operatorname{H}^{2}(p_{i},p_{j})\Big).$
We also need the following result.
Lemma 11 (Lemma 32.11 of [50]).
For any $p\in\mathcal{P}$, $\eta\geqslant\varepsilon$, and an integer $k\geqslant 0$,
$N_{\operatorname{H}}(B_{\operatorname{H}}(p,2^{k}\eta)\cap\mathcal{P},\eta/2)\leqslant N_{\mathsf{loc}}(\mathcal{P},\varepsilon)^{k+1}.$
Proof of Lemma˜9.
Set $\varepsilon=\varepsilon_{n}$ and $R=4\varepsilon$. Let $\{p_{1},\ldots,p_{N}\}$ be a maximal $\varepsilon$-packing. There is an index, which we rename as $1$, such that $\operatorname{H}(p^{\star},p_{1})\leqslant\varepsilon$. We run the tests, compute $T_{i}$, and choose $\widehat{p}=p_{i^{\star}}$ as above. By the triangle inequality and the definition of $i^{\star}$,
$\operatorname{H}(\widehat{p},p^{\star})\leqslant\operatorname{H}(\widehat{p},p_{1})+\operatorname{H}(p_{1},p^{\star})\leqslant\max\{R,T_{1}\}+\varepsilon\leqslant 5\varepsilon+T_{1}.$
Hence, for any $t\geqslant 4$,
$\mathbf{P}\big(\operatorname{H}(\widehat{p},p^{\star})>5\varepsilon+4t\varepsilon\big)\leqslant\mathbf{P}(T_{1}\geqslant tR).$
Define the Hellinger shells $A_{k}=\{p\in\mathcal{P}:2^{k}R\leqslant\operatorname{H}(p_{1},p)<2^{k+1}R\}$ and set $G_{k}=\{p_{1},\ldots,p_{N}\}\cap A_{k}$. By a union bound over $G_{k}$ and Lemma 10,
$\mathbf{P}(T_{1}\in[2^{k}R,2^{k+1}R))\leqslant\sum_{j:\ p_{j}\in G_{k}}\exp\Big(-\frac{n}{4}\operatorname{H}^{2}(p_{1},p_{j})\Big)\leqslant|G_{k}|\exp\Big(-\frac{n}{4}(2^{k}R)^{2}\Big).$
Since $\operatorname{H}(p_{1},p^{\star})\leqslant\varepsilon$, we have
$B_{\operatorname{H}}(p_{1},2^{k+1}R)\subset B_{\operatorname{H}}(p^{\star},2^{k+1}R+\varepsilon)\subset B_{\operatorname{H}}(p^{\star},2^{k+4}\varepsilon),$
and by Lemma 11,
$|G_{k}|\leqslant N_{\operatorname{H}}\big(B_{\operatorname{H}}(p_{1},2^{k+1}R)\cap\mathcal{P},\ \varepsilon/2\big)\leqslant N_{\mathsf{loc}}(\mathcal{P},\varepsilon)^{k+5}.$
Using the fixed point,
$\mathbf{P}(T_{1}\geqslant tR)\leqslant\sum_{k\geqslant\lfloor\log_{2}t\rfloor}\exp\big(n\varepsilon^{2}(k+5)-4n\varepsilon^{2}\,4^{k}\big)\leqslant\sum_{k\geqslant\lfloor\log_{2}t\rfloor}\exp\big(-n\varepsilon^{2}\,4^{k}\big).$
Since $2^{\lfloor\log_{2}t\rfloor}\leqslant t<2^{\lfloor\log_{2}t\rfloor+1}$ implies $4^{\lfloor\log_{2}t\rfloor}\geqslant t^{2}/4$, we have $\mathbf{P}(T_{1}\geqslant tR)\leqslant 2\,\exp\Big(-\frac{n\varepsilon^{2}\,t^{2}}{4}\Big).$ We choose $t=\max\Big\{4,\sqrt{\frac{4}{n\varepsilon^{2}}\log\frac{2}{\delta}}\ \Big\},$ which implies $2\exp(-n\varepsilon^{2}t^{2}/4)\leqslant\delta$. Therefore, with probability at least $1-\delta$,
$\operatorname{H}(\widehat{p},p^{\star})\leqslant 5\varepsilon+4t\varepsilon\leqslant 21\varepsilon+8\sqrt{\frac{\log(2/\delta)}{n}}.$
This gives the claim. Finally, for any $s\geqslant 0$, choosing $\delta(s)=\min\left\{1,2\exp(-ns^{2}/64)\right\}$ gives
$\mathbf{P}\left(\operatorname{H}(p^{\star},\widehat{p})>21\varepsilon_{n}+s\right)\leqslant\min\left\{1,2\exp(-ns^{2}/64)\right\}.$
Let $s_{0}=8\sqrt{\log 2/n}$. Integrating the tail, we have
${\mathbf{E}}\operatorname{H}(p^{\star},\widehat{p})\leqslant 21\varepsilon_{n}+\int_{0}^{s_{0}}1\,ds+\int_{s_{0}}^{\infty}2\exp(-ns^{2}/64)ds\leqslant 21\varepsilon_{n}+\frac{11}{\sqrt{n}}.$
The second claim follows. ∎
Appendix D Other deferred proofs
D.1 Proof of Lemma˜1
Let h(x)=p(x)/q(x)h(x)={p(x)}/{q(x)}. By the standard representation we have
$\operatorname{KL}(p,q)=\int p\log\frac{p}{q}\,d\mu=\int q\bigl(h\log h-h+1\bigr)\,d\mu.$
We invoke the elementary pointwise inequality, valid for all $h>0$,
$h\log h-h+1\leqslant\frac{1}{(\sqrt{e}-1)^{2}}(\sqrt{h}-1)^{2}\max\{1,\log h\}.$ (31)
The inequality (31) can be shown by treating $h\leqslant e$ and $h>e$ separately and checking that $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}}$ is increasing on $(0,e]$ while $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}\log h}$ is decreasing on $e,\infty)$; both meet at value $(\sqrt{e}-1)^{-2}$ when $h=e$. Applying (31) with $h=p/q$ and $m=\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\right\}$, we obtain
$\operatorname{KL}(p,q)\leqslant\frac{m}{(\sqrt{e}-1)^{2}}\int q(\sqrt{h}-1)^{2}d\mu.$
Finally, observe that $\int q(\sqrt{h}-1)^{2}\,d\mu=\int(\sqrt{p}-\sqrt{q})^{2}d\mu=2\operatorname{H}^{2}(p,q)$. Substituting yields the claim. ∎
D.2 Proof of Corollary˜2
We start with a general amplification lemma that converts a constant-factor ratio cover into a $(1+\varepsilon)$-ratio cover.
Claim 2.
For $0<\varepsilon\leqslant 1$ and $C\geqslant 1$, let $B\subset K$ be a $C$-ratio cover of a convex and compact set $K\subset\mathbb{R}_{+}^{d}$. Then there exists a subset $A\subset K$ with
$|A|\leqslant|B|\left(3+\frac{4\log_{2}(4C/\varepsilon)}{\varepsilon}\right)^{d-1}$
that is a $(1+\varepsilon)$-ratio cover of $K$.
Proof.
Fix $b\in B$. We will add to $A$ several points of the form
$z=(\varepsilon/4)b+(1-\varepsilon/4)w,$
where $w\in K$ is chosen so that its first d1d-1 coordinates fall into prescribed multiplicative bins relative to $b$.
We now describe this discretization. Fix $\theta\in K$, and choose $b\in B$ that $C$-ratio covers $\theta$ (if there are several, choose one arbitrarily). Then for each coordinate $i\in[d]$ we have $\theta_{i}\leqslant Cb_{i}$, and hence
$\theta_{i}\leqslant(\varepsilon/4)b_{i}\cdot\frac{4C}{\varepsilon}.$ (32)
We split the possible values of $\theta_{i}$ into the following cases relative to $b_{i}$:
  • Case 1: $\theta_{i}\leqslant(\varepsilon/4)\,b_{i}$;
  • Case 2: $\theta_{i}>(\varepsilon/4)\,b_{i}$, in which case (32) implies that
    $\theta_{i}\in\Bigl[(\varepsilon/4)\,b_{i}\,(1+\varepsilon/4)^{j},\ (\varepsilon/4)\,b_{i}\,(1+\varepsilon/4)^{j+1}\Bigr]\quad\text{for some }j\in\Bigl\{0,\dots,\Bigl\lceil\log_{1+\varepsilon/4}\Bigl(\frac{4C}{\varepsilon}\Bigr)\Bigr\rceil\Bigr\}.$
We now define the set $A$. For each $b\in B$, and for each possible choice of cases/bins for the first d1d-1 coordinates, we include one point
$z=(\varepsilon/4)\,b+(1-\varepsilon/4)\,w,$
where $w\in K$ has the specified cases/bins on coordinates $1,\dots,d-1$ (if such a $w$ exists). If such a $w$ exists, we choose $w$ to maximize the $d$-th coordinate among all points in $K$ with this bin pattern. This maximizer exists since $K$ is compact and the bin constraints above are closed. There are at most $\bigl(2+\lceil\log_{1+\varepsilon/4}(4C/\varepsilon)\rceil\bigr)^{d-1}$ bin patterns for the first d1d-1 coordinates, so
$|A|$ $\leqslant|B|\Bigl(2+\Bigl\lceil\log_{1+\varepsilon/4}\Bigl(\frac{4C}{\varepsilon}\Bigr)\Bigr\rceil\Bigr)^{d-1}\leqslant|B|\Bigl(3+\frac{\log_{2}(4C/\varepsilon)}{\log_{2}(1+\varepsilon/4)}\Bigr)^{d-1}$
$\leqslant|B|\Bigl(3+\frac{4\log_{2}(4C/\varepsilon)}{\varepsilon}\Bigr)^{d-1},$
where in the last step we used $\log_{2}(1+x)\geqslant x$ for $0\leqslant x\leqslant 1$ and $\varepsilon/4\leqslant 1/4$.
It remains to show that $A$ is a $(1+\varepsilon)$-ratio cover of $K$. Fix $\theta\in K$, and let $b\in B$ be a $C$-ratio cover of $\theta$. By construction of $A$, there exists $z=(\varepsilon/4)\,b+(1-\varepsilon/4)\,w\in A$ such that $w$ has the same bin pattern as $\theta$ on the first d1d-1 coordinates, and in addition $\theta_{d}\leqslant w_{d}$ (since $w$ maximizes the $d$-th coordinate within that pattern). We check coordinatewise domination.
For $i\in[d-1]$, if $\theta_{i}$ falls in Case 1 then $\theta_{i}\leqslant(\varepsilon/4)\,b_{i}\leqslant z_{i}$. Otherwise $\theta_{i}$ is in Case 2, and $\theta_{i}$ and $w_{i}$ lie in the same interval $[L,U]$ with $U=(1+\varepsilon/4)L$ and $w_{i}\geqslant L$, so $\theta_{i}\leqslant U\leqslant(1+\varepsilon/4)\,w_{i}$. Using $z_{i}\geqslant(1-\varepsilon/4)\,w_{i}$,
$\theta_{i}\leqslant(1+\varepsilon/4)\,w_{i}\leqslant(1+\varepsilon/4)\,\frac{z_{i}}{1-\varepsilon/4}\leqslant(1+\varepsilon/4)(1+\varepsilon/2)\,z_{i}\leqslant(1+\varepsilon)\,z_{i},$
where we used $1/(1-x)\leqslant 1+2x$ for $x\in[0,1/2]$ (here $x=\varepsilon/4$) and the fact that $(1+\varepsilon/4)(1+\varepsilon/2)\leqslant 1+\varepsilon$ for $0<\varepsilon\leqslant 1$. Finally, for the last coordinate, $\theta_{d}\leqslant w_{d}\leqslant z_{d}/(1-\varepsilon/4)\leqslant(1+\varepsilon)\,z_{d}$. Thus $\theta_{i}\leqslant(1+\varepsilon)z_{i}$ holds for all $i\in[d]$, proving that $A$ is a $(1+\varepsilon)$-ratio cover. ∎
To conclude the theorem, apply ˜2 with $K$ as given, $C=32$, and $B$ a $32$-ratio cover of $K$ of size at most $2^{8d}$ provided by Theorem˜3. This yields an $\varepsilon$-approximate Pareto curve of size
$|A|\leqslant 2^{8d}\,\Bigl(3+\frac{4\log_{2}(4\cdot 32/\varepsilon)}{\varepsilon}\Bigr)^{d-1}=2^{8d}\,\Bigl(3+\frac{4\log_{2}(128/\varepsilon)}{\varepsilon}\Bigr)^{d-1}.$
The claim follows. ∎
D.3 Proof of Lemma˜5
Fix $q\in\mathcal{P}$ and set $\Delta=1\vee\operatorname{H}^{2}(p,q)/\varepsilon$ and $\lambda=1/\Delta$. Define $q_{\lambda}=(1-\lambda)p+\lambda q\in\mathcal{P}$. By convexity of $\operatorname{H}^{2}(p,\cdot)$ we have
$\operatorname{H}^{2}(p,q_{\lambda})=\operatorname{H}^{2}\bigl(p,(1-\lambda)p+\lambda q\bigr)\leqslant\lambda\,\operatorname{H}^{2}(p,q)\leqslant\varepsilon,$
so $q_{\lambda}\in\mathcal{Q}$. Pick $\widetilde{q}\in S$ such that $q_{\lambda}\leqslant\alpha\,\widetilde{q}$ pointwise. Since $q_{\lambda}\geqslant\lambda q$, it follows that $q\leqslant\alpha\Delta\,\widetilde{q}$ and hence
$\sup_{x\in\mathcal{X}}\log\Bigl(\frac{q(x)}{\widetilde{q}(x)}\Bigr)\leqslant\log(\alpha\Delta).$
Also $\widetilde{q}\in\mathcal{Q}$, so $\operatorname{H}^{2}(p,\widetilde{q})\leqslant\varepsilon$, and by the triangle inequality, it holds that
$\operatorname{H}^{2}(q,\widetilde{q})=\operatorname{H}(q,\widetilde{q})^{2}\leqslant\bigl(\operatorname{H}(q,p)+\operatorname{H}(p,\widetilde{q})\bigr)^{2}\leqslant 2\operatorname{H}^{2}(p,q)+2\operatorname{H}^{2}(p,\widetilde{q})\leqslant 2\bigl(\operatorname{H}^{2}(p,q)+\varepsilon\bigr).$
Applying Lemma 1 gives
[…]
$\leqslant\frac{4}{(\sqrt{e}-1)^{2}}(\operatorname{H}^{2}(p,q)+\varepsilon)\max\{1,\log(\alpha\Delta)\}.$
Since $\alpha\geqslant 2$ and $\Delta\geqslant 1$, we have $\log(\alpha\Delta)\geqslant\log 2$, so $\max\{1,\log(\alpha\Delta)\}\leqslant(\log(\alpha\Delta))/\log 2$. Using $\frac{4}{(\sqrt{e}-1)^{2}\log 2}<30$, we obtain
$\operatorname{KL}(q,\widetilde{q})\leqslant 30(\operatorname{H}^{2}(p,q)+\varepsilon)\log(\alpha\Delta).$
Finally, $\log(\alpha\Delta)=\max\left\{\log(\alpha),\ \log\left(\alpha\operatorname{H}^{2}(p,q)/\varepsilon\right)\right\}$ by definition of $\Delta$. Taking the minimum over $\widetilde{q}\in S$ yields the claim.∎
D.4 Proof of Proposition˜3
First, we present the required estimator, which is simpler than the one we used in Theorem˜1.
  • Construct a class of densities
    $\mathcal{Q}=\left\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\mathcal{P}\right\}.$ (33)
  • Run the density estimator of Lemma˜2 over $\mathcal{Q}$ to output the final density $\widehat{p}$.
Let $p^{\star}_{\mathcal{P}}$ be any minimizer of $\operatorname{KL}(p^{\star},p)$ over $p\in\mathcal{P}$ and let
$\widetilde{p}=\left(1-\frac{1}{n}\right)p^{\star}_{\mathcal{P}}+\frac{1}{nM}\sum_{q\in\mathcal{P}}q.$
Observe that $\widetilde{p}\in\mathcal{Q}$ and that for each $p,q\in\mathcal{Q}$ we have
$\sup_{p,q\in\mathcal{Q}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+nM\bigr)\leqslant\log(2nM).$
Therefore, by Lemma˜2 we have, with probability at least $1-\delta$,
${\mathbf{E}}\log(\widehat{p}(X))+{\mathbf{E}}\log(\widetilde{p}(X))\leqslant\frac{2\log(M)+\frac{25}{4}\left(e-2\right)\left(\log(2nM)+1\right)\log\left(\frac{1}{\delta}\right)}{n}.$
Finally, we show that
${\mathbf{E}}\log\bigl(p^{\star}_{\mathcal{P}}(X)\bigr)\leqslant{\mathbf{E}}\log\bigl(\widetilde{p}(X)\bigr)+\frac{2}{n},$
which, by combining this with the previous inequality and simplifying the constants, immediately proves the claim. Indeed, by the definition of $\widetilde{p}$ we have, for all $x\in\mathcal{X}$,
$\log\left(\frac{p^{\star}_{\mathcal{P}}(x)}{\widetilde{p}(x)}\right)\leqslant\log\left(\frac{1}{1-1/n}\right)\leqslant\frac{2}{n},$ (34)
whenever $n\geqslant 2$. When $n=1$, we have $\mathcal{Q}=\big\{\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q\big\}$ and the claim follows since $\widehat{p}=\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q$.∎
D.5 Proof of Proposition˜4
We assume throughout that $n\geqslant 2$. When $n=1$, $\mathcal{Q}$ contains a single element, namely $\mathcal{Q}=\left\{\frac{1}{M}\sum_{q\in\mathcal{P}}q\right\},$ so $\widehat{p}$ is this deterministic density and the stated bound is immediate.
The proof exploits the curvature of the logarithmic loss on the ratio-bounded convex class $\mathcal{Q}$ via an offset-process argument (see [42, 44, 58] for related derivations). Recall that $\widehat{p}$ is the maximum-likelihood estimator over $\mathcal{Q}$, namely
$\widehat{p}\in\arg\min_{p\in\mathcal{Q}}\widehat{L}(p),\qquad\widehat{L}(p)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}),\qquad L(p)=-{\mathbf{E}}\log p(X),$
where $X\sim p^{\star}$ and $\mathbf{E}$} denotes expectation under $p^{\star}$. As in the proof of Proposition˜3, the smoothing in the definition of $\mathcal{Q}$ implies that
$\sup_{p,q\in\mathcal{Q}}\sup_{x\in\mathcal{X}}\Bigl|\log\Bigl(\frac{p(x)}{q(x)}\Bigr)\Bigr|\leqslant\log(2nM).$ (35)
We start with the following symmetrization lemma.
Lemma 12.
Let $p^{\star}_{\mathcal{Q}}\in\arg\min_{p\in\mathcal{Q}}L(p)$. For every $\lambda>0$ it holds that
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}Z(p)\Big),$ (36)
where
$Z(p)=\frac{1}{n}\sum_{i=1}^{n}\left[4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right],$
and $(\varepsilon_{i})_{i=1}^{n}$ are i.i.d. Rademacher signs independent of $(X_{i})_{i=1}^{n}$.
Proof.
Define the log-density ratio $r_{p}(x)=\log\left(\frac{p^{\star}_{\mathcal{Q}}(x)}{p(x)}\right).$ Let $(X_{1}^{\prime},\dots,X_{n}^{\prime})$ be an independent ghost sample (i.i.d. as $(X_{i})$) and set
$P_{n}r_{p}=\frac{1}{n}\sum_{i=1}^{n}r_{p}(X_{i}),\quad P_{n}(r_{p}^{2})=\frac{1}{n}\sum_{i=1}^{n}r_{p}(X_{i})^{2},\quad P_{n}^{\prime}r_{p}=\frac{1}{n}\sum_{i=1}^{n}r_{p}(X_{i}^{\prime}),\quad P_{n}^{\prime}(r_{p}^{2})=\frac{1}{n}\sum_{i=1}^{n}r_{p}(X_{i}^{\prime})^{2}.$
By convexity of $\mathcal{Q}$, the midpoint $\bar{p}=\frac{\widehat{p}+p^{\star}_{\mathcal{Q}}}{2}$ belongs to $\mathcal{Q}$. By the definition of $\widehat{p}$, it holds that
$\widehat{L}(\widehat{p})\leqslant\widehat{L}(\bar{p}).$ (37)
Using (35) we can apply Lemma˜7, which gives for all $i\in[n]$,
$\log(\bar{p}(X_{i}))=-\log\Big(\tfrac{\widehat{p}(X_{i})+p^{\star}_{\mathcal{Q}}(X_{i})}{2}\Big)\leqslant\frac{-\log\widehat{p}(X_{i})-\log p^{\star}_{\mathcal{Q}}(X_{i})}{2}-\frac{1}{8(\log(2nM)+1)}\,r_{\widehat{p}}(X_{i})^{2}.$
Averaging over $i\in[n]$, combining with (37) and rearranging we obtain
$P_{n}r_{\widehat{p}}\leqslant-\frac{1}{4(\log(2nM)+1)}P_{n}\big(r_{\widehat{p}}^{2}\big).$ (38)
We repeat the same lines at the population level. Fix any $p\in\mathcal{Q}$ and let $\widetilde{p}=(p+p^{\star}_{\mathcal{Q}})/2\in\mathcal{Q}$. Because $p^{\star}_{\mathcal{Q}}$ is a population minimizer over $\mathcal{Q}$, we have
[…]
Applying Lemma 7 as above, using (39), taking the expectation and rearranging we obtain that for all $p\in\mathcal{Q}$,
${\mathbf{E}}[r_{p}(X)]\geqslant\frac{1}{4(\log(2nM)+1)}{\mathbf{E}}[r_{p}(X)^{2}].$ (40)
We have $L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})={\mathbf{E}}r_{\widehat{p}}(X)$. For brevity of notation, set $\alpha=\frac{1}{4(\log(2nM)+1)}.$ The following lines of inequalities hold (conditioned on $\widehat{p}$):
[…]
$=2\Big(\big({\mathbf{E}}-P_{n}\big)r_{\widehat{p}}+P_{n}r_{\widehat{p}}-\frac{\alpha}{2}{\mathbf{E}}\big[r_{\widehat{p}}(X)^{2}\big]\Big)$
[…]
$\leqslant\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-2\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}$
$\leqslant\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha P_{n}\big(r_{p}^{2}\big)-\alpha{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}.$
Finally, combining the above inequality with the standard symmetrization argument (that is, when appropriate $\mathbf{E}$} corresponds to taking the expectation with respect to the joint distribution of $(X_{i})$, $(X_{i}^{\prime})$ and $(\varepsilon_{i})$) we have for every $\lambda>0$,
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)$ $\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}\big)$
$\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{Q}}\Big\{2\big(P_{n}^{\prime}-P_{n}\big)r_{p}-{\alpha}P_{n}\big(r_{p}^{2}\big)-{\alpha}P_{n}^{\prime}\big(r_{p}^{2}\big)\Big\}\big),$
$={\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}\frac{1}{n}\sum_{i=1}^{n}\Big[2\varepsilon_{i}\big(r_{p}(X_{i}^{\prime})-r_{p}(X_{i})\big)-\alpha\big(r_{p}(X_{i})^{2}+r_{p}(X_{i}^{\prime})^{2}\big)\Big]\Big)$
$\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}\frac{1}{n}\sum_{i=1}^{n}\Big[4\varepsilon_{i}r_{p}(X_{i})-2\alpha r_{p}(X_{i})^{2}\Big]\Big),$
where in the last line we split the terms involving $P_{n}$ and $P_{n}\prime$ and use that for a pair of random variables $Y,Y^{\prime}$ having the same distribution, it holds that ${\mathbf{E}}\exp(Y+Y^{\prime})\leqslant\sqrt{{\mathbf{E}}\exp(2Y)}\sqrt{{\mathbf{E}}\exp(2Y^{\prime})}={\mathbf{E}}\exp(2Y)$. The claim follows. ∎
The next lemma is about the Lipschitzness of the loss class indexed by $\mathcal{Q}$. For $\alpha\in\Delta_{M-1}$, we denote
$w_{\alpha}(x)=\Bigl(1-\frac{1}{n}\Bigr)\sum_{j=1}^{M}\alpha_{j}p_{j}(x)+\frac{1}{nM}\sum_{j=1}^{M}p_{j}(x),$
so that $\mathcal{Q}=\{w_{\alpha}:\alpha\in\Delta_{M-1}\}$, where $\Delta_{M-1}$ is the simplex in $\mathbb{R}^{M}$.
Lemma 13.
For any $\alpha,\beta\in\Delta_{M-1}$ it holds that
$\sup\limits_{x\in\mathcal{X}}|\log(w_{\alpha}(x))-\log(w_{\beta}(x))|\leqslant nM\|\alpha-\beta\|_{1}.$
Proof.
Fix $\alpha,\beta\in\Delta_{M-1}$ and $x\in\mathcal{X}$. It holds that
$|w_{\alpha}(x)-w_{\beta}(x)|=\left(1-\frac{1}{n}\right)\left|\sum\limits_{j=1}^{M}(\alpha_{j}-\beta_{j})p_{j}(x)\right|\leqslant\sum_{j=1}^{M}|\alpha_{j}-\beta_{j}|\,p_{j}(x)\leqslant\|\alpha-\beta\|_{1}\sum_{j=1}^{M}p_{j}(x).$
By the mean value theorem we have
$|\log(w_{\alpha}(x))-\log(w_{\beta}(x))|\leqslant\frac{|w_{\alpha}(x)-w_{\beta}(x)|}{\min\{w_{\alpha}(x),w_{\beta}(x)\}}.$
Observe that $w_{\alpha}(x)\geqslant\frac{1}{nM}\sum_{j=1}^{M}p_{j}(x)$ and $w_{\beta}(x)\geqslant\frac{1}{nM}\sum_{j=1}^{M}p_{j}(x)$. Combining the above inequalities yields
$|\log(w_{\alpha}(x))-\log(w_{\beta}(x))|\leqslant\frac{\|\alpha-\beta\|_{1}\sum_{j=1}^{M}p_{j}(x)}{\frac{1}{nM}\sum_{j=1}^{M}p_{j}(x)}=nM\|\alpha-\beta\|_{1}.$
The claim follows. ∎
We fix $\rho\in(0,1]$ and let $\mathcal{A}_{\rho}$ be the minimal $\rho$-net of $\Delta_{M-1}$ with respect to the $\|\cdot\|_{1}$ distance. Define the finite subclass $\mathcal{Q}_{\rho}=\{w_{\alpha}:\alpha\in\mathcal{A}_{\rho}\}\subseteq\mathcal{Q}$. Let $p$ be any element in $\mathcal{Q}$ and $p_{\rho}$ be the $\rho$-close element to it in $\mathcal{Q}_{\rho}$ with respect to the net $\mathcal{A}_{\rho}$. By Lemma˜13 and (35) we have
$Z(p)-Z(p_{\rho})\leqslant 4nM\rho+\frac{\log(2nM)}{\log(2nM)+1}nM\rho\leqslant 5nM\rho,$
which implies for any $\lambda>0$,
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}Z(p)\Big)$ $\leqslant\exp\Big(5\lambda nM\rho\Big)\,{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\max_{q\in\mathcal{Q}_{\rho}}Z(q)\Big)$
$\leqslant|\mathcal{Q}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\max\limits_{p\in\mathcal{Q}_{\rho}}{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big).$
Using ${\mathbf{E}}\exp(\lambda\varepsilon_{i})\leqslant\exp(\lambda^{2}/2)$ we have for any $\lambda\in\left[0,\frac{n}{16(\log(2nM)+1)}\right]$ and $p\in\mathcal{Q}$,
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big)$ $={\mathbf{E}}_{\varepsilon}\exp\left(\frac{\lambda}{n}\sum\limits_{i=1}^{n}\left(4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right)\right)$
$\leqslant\exp\left(\sum\limits_{i=1}^{n}\left(\frac{8\lambda^{2}}{n^{2}}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{\lambda}{2n(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right)\right)$
$\leqslant 1.$
Taking expectation over $(X_{i})$ and using (36), we obtain that for any $\lambda\in\left[0,\frac{n}{16(\log(2nM)+1)}\right]$,
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)\leqslant|\mathcal{Q}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\leqslant\exp\left(M\log(3/\rho)+5\lambda nM\rho\right),$ (41)
where we used the standard volumetric argument (see [61, Lemma 5.7]) to bound $|\mathcal{Q}_{\rho}|$ for $\rho\leqslant 1$. This immediately implies that, with probability at least $1-\delta$,
$L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\leqslant 5nM\rho+\frac{1}{\lambda}\Big(M\log(3/\rho)+\log(1/\delta)\Big).$ (42)
Taking $\lambda=\frac{n}{16(\log(2nM)+1)}$ and $\rho=\frac{1}{3n^{2}M}\leqslant 1$, we obtain, using $L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})=\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})$,
$\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})\leqslant\frac{5}{3n}+\frac{16(\log(2nM)+1)(M\log(9n^{2}M)+\log(1/\delta))}{n}.$ (43)
Finally, repeating the argument in (34) we have
$\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})\leqslant\inf_{p\in\mathop{\mathrm{conv}}(\mathcal{P})}\operatorname{KL}(p^{\star},p)+\frac{2}{n}.$ (44)
The claim follows by combining (43) and (44) and adjusting the constants. ∎